### v1

# Designing Elegant Bell Inequalities

Abstract
Elegant Bell inequality is well known for its much exploited property, being maximally violated by maximal entanglement, mutually unbiased bases, and symmetric informationally complete positive operator-valued measure elements. It is the only one with such property known so far. We present a method to construct Bell inequalities with violation feature analogous to original elegant Bell inequality in high dimension from a simple analytic quantum bound. A Bell inequality with such feature is derived in three dimension for the first time. It shows larger violation than existing Bell inequalities of similar classes while requiring arguably small number of measurements.
I Introduction
Performing local measurements on the shared system, distant subsystems can reveal a quantum correlation that cannot be reproduced by any local hidden variable theoretical explanation. Such nonlocality of the correlation can be tested via the violation of Bell inequalities [1, 2]. What Bell inequality can tell us about the system goes beyond the scope of the mere existence of nonlocal correlation. It has been investigated the method to exploit Bell’s theorem for further certifying the properties of given system in so-called device-independent (DI) approach [3, 4]. This approach leads to various applications, including DI quantum key distribution [4, 5, 6, 7, 8] and DI certification of randomness [9, 10, 11].
Reflecting on the aforementioned possibilities provided by Bell inequalities, it is of prime importance to derive them such that their maximal violations are obtained from desirable features, for example maximal entanglement [12, 13, 14, 15, 16, 17, 18, 19, 20], mutually unbiased bases (MUBs) [19, 20], or symmetric informationally complete POVM elements (SICs) [21, 20]. As is well known, Clauser-Horne-Shimony-Holt Bell inequality is the case in which the maximal violation can be obtained from maximal entanglement and Pauli operators having MUBs as their eigenstates. There exists another bipartite Bell inequality so-called Gisin’s elegant Bell inequality (EBI) [21],
$E_{11}+E_{12}-E_{13}-E_{14}$
$+E_{21}-E_{22}+E_{23}-E_{24}$
$+E_{31}-E_{32}-E_{33}+E_{34}\leq 6,$
also optimally violated by maximally entangled state. The expectation value of the product outcomes of Alice’s $x$-th measurement and Bob’s $y$-th measurement is denoted by $E_{xy}$. Quantum bound for EBI is proven as $64\sqrt{3}\simeq 6.9282>$ 6 [22]. What makes EBI more interesting is that its violation is obtained from the case where the local measurements of each subsystem is constructed by MUBs and SICs respectively. As first remarked by Gisin, the geometrical symmetries in the optimal measurement settings of EBI are clearly revealed when they are mapped on Bloch’s sphere [21]. Three eigenstates defining three projective measurements of Alice form three mutually orthogonal vectors on Bloch sphere and four eigenstates of Bob’s side make a tetrahedron [21]. Original EBI has been used in several fields of quantum information, including DI randomness certification [22, 23], DI certification of non-projective measurement [24, 25], and for other tasks [26, 27, 28]. Self-testing property of EBI has also been studied in [29].
MUBs and SICs are important resources in quantum cryptography [30, 31, 32, 22, 23]. There has been efforts to derive Bell inequalities for MUBs or SICs exploiting the advantage of high dimension [19, 20]. For DI randomness certification, for which high dimension provides natural advantage [22], Bell inequalities for SICs has been used [22, 23, 20]. EBI is distinguished from ones either for MUBs or SICs as it is optimally violated by both of them.
We address the problem of deriving Bell inequalities with violation feature analogous to original EBI in high dimensions. The problem is closely related to the question of what benefit for Bell nonlocality would come from simultaneous consideration of the symmetries of MUBs and SICs. It is worth noting that although EBI is originally derived in the different context [33, 21], we focus on the generalization of the property, being maximally violated by maximal entanglement, MUBs and SICs. Such generalization of EBI has not much been investigated so far to our knowledge. We approach the raised problem by defining a generalized type of correlation as follows.
Definition. For prime $d\geq 3$, generalized EBI correlation is defined as the bipartite correlation obtained from the quantum realization: (i) maximally entangled state $\left|\phi_{d}^{+}\right>:=(1/\sqrt{d})\sum_{\alpha=0}^{d-1}\left|\alpha\alpha\right>$, (ii) Alice’s $d^{2}-1$ observables comprising $\mathcal{W}_{d}\setminus\mathds{1}$ and (iii) Bob’s $d^{2}$ observables having $d^{2}$ SICs in their eigenspaces.
In condition (ii), $\mathcal{W}_{d}:=\{X^{m}Z^{n}|m,n\in[0,d)\}$ is the set of $d^{2}$ Weyl-Heisenberg (WH) operators defined with $X:=\sum_{\alpha=0}^{d-1}\left|\alpha+1\right>\left<\alpha\right|$, $Z:=\sum_{\alpha=0}^{d-1}\omega^{\alpha}\left|\alpha\right>\left<\alpha\right|$ and $\omega:=e^{2\pi i/d}$. The integer domain $\{0,1,\ldots,d-1\}$ is denoted by $[0,d)$ and similar abbreviation is to be used throughout this work. Note that $X^{0}Z^{0}=\mathds{1}$ is not included in the condition (ii). Conditions (i) and (iii) are natural generalization of optimal quantum realization of original EBI. As for the generalization of Alice’s optimal observables, two interpretations are possible. The set of Alice’s optimal observables for EBI, $\sigma_{x},\sigma_{y},\sigma_{z}$, [22] can be viewed either as “$d+1$ observables” defining $d+1$ sets of MUBs, or as “$d^{2}-1$ unitary operator bases” [34] for a traceless operator $\vec{a}\cdot\vec{\sigma}$ defined with a three-dimensional unit vector $\vec{a}$ and $\vec{\sigma}:=(\sigma_{x},\sigma_{y},\sigma_{z})$ when $d=2$. The above duality arises as $d+1=d^{2}-1$ when $d=2$. Condition (ii) clearly generalizes the latter. Under condition (ii), complete MUBs are realized in eigenspace of $d+1$ among $d^{2}-1$ observables of Alice, namely $Z$, $X$, $XZ$, …, $XZ^{d-1}$ [35]. Moreover, the eigenbases of all $d^{2}-1$ observables of Alice comprises the $d-1$ same sets of complete $d+1$ MUBs. It can be easily seen from the fact that $XZ^{n}$ and $X^{m}Z^{mn}$ has the same eigenbases for all $m$ and $\mathcal{W}=\{X^{m}Z^{mn}|m,n\in[0,d)\}$ for prime $d$. In the following, we will further explain the reason for the choice of optimal observables of Alice as unitary operator bases.
Figure 1: On the shared system, two subsystems, Alice and Bob, respectively performs one of $d^{2}-1$ and $d^{2}$ measurements outputting one of $d$ outcomes.
Another insight that underlies our approach is to exploit $systematic$ and $versatile$ features of WH group covariant SICs. For brevity, consider the parameterization $W_{dp+q}:=X^{p}Z^{q}$, such that $\mathcal{W}_{d}=\{W_{j}|j\in[0,d^{2})\}$. Group covariant SICs is derived by acting group elements on a single state, so-called fiducial vector [36]. For the case of WH group covariant SICs, group elements $W_{j}$ generates $d^{2}$ SICs on the support of fiducial state $\left|\varphi\right>$, $\{W_{j}\left|\varphi\right>|j\in[0,d^{2})\}$. Weyl-Heisenberg group covariance plays crucial role for satisfying the condition (iii) and provides systematic feature for our formalism. Versatility of our method is from the fact that it can be applied only if WH group covariant SICs for given local dimension exists. The existence of SICs are known in all dimension up to $d=193$ [37]. More importantly, all group covariant SICs is covariant to WH group in prime dimensions [38].
II Preliminaries
We consider the measurement scenario of two subsystems Alice and Bob respectively performing one of possible $d^{2}-1$ and $d^{2}$ number of $d$ outcome measurements $A_{x}$ and $B_{y}$ on the shared system. Measurements are labelled with $x\in\{1,2,\ldots,d^{2}-1\}$ and $y\in\{0,1,\ldots,d^{2}-1\}$. Outcomes of $A_{x}$ and $B_{y}$ are respectively distinguished with indices $\alpha,\beta\in\{0,1,\ldots,d-1\}$. The scenario is illustrated in Fig. 1. Probability of obtaining the outcomes $\alpha$ and $\beta$ respectively from settings $x$ and $y$ is denoted by $P(\alpha\beta|xy)$. A statistics of the Bell test can be described with correlation, $\mathbf{p}:=\{P(\alpha\beta|xy)|\forall\alpha,\beta,x,y\}$. Quantum correlation allows
$P(\alpha\beta|xy)=\operatorname{tr}\left[\rho_{AB}(G_{\alpha}^{x}\otimes H_{\beta}^{y})\right]$
where $\rho_{AB}$ is the bipartite density matrix of Alice and Bob. Measurement operators, $\{G_{\alpha}^{x}\}$ and $\{H_{\beta}^{y}\}$, define the measurements of Alice and Bob respectively. Bell inequality is used to show that there is quantum correlation which does not allow the local hidden variable (LHV) theoretical description of the correlation [1]. Geometrically, LHV correlation region forms a polytope, $\mathcal{L}$, and quantum region, $\mathcal{Q}$, is known to include it [2].
The generic form of so-called Bell expression [2] without marginal probabilities can be defined as,
$S$ $=\sum_{\alpha,\beta}\sum_{x,y}g_{x,y}^{\alpha,\beta}P(\alpha\beta|xy)$ (1)
$=\sum_{n}\sum_{x,y}f_{x,y}^{n}\langle A_{x}^{n}B_{y}^{n}\rangle.$ (2)
where $n\in(1,d)$ and $g_{x,y}^{\alpha,\beta}\in\mathbb{R}$ is the real weight for each probability. For simplicity, the summations over all possible values are to be denoted without limits. The correlator $\langle A_{x}^{n}B_{y}^{n}\rangle=\sum_{\alpha,\beta=0}^{d-1}\omega^{n(\alpha+\beta)}P(\alpha\beta|xy)$, where the outcomes of the measurements $A_{x}$ and $B_{y}$ are set as $\omega^{\alpha},\omega^{\beta}\in\{1,\omega,\omega^{2},\ldots,\omega^{d-1}\}$. And the coefficient $g_{x,y}^{\alpha,\beta}:=\sum_{n}f_{x,y}^{n}\omega^{n(\alpha+\beta)}$ The condition $f_{x,y}^{d-n}:=(f_{x,y}^{n})^{*}$ guarantees the real-valuedness of $S$. A expression is specified by a $(d^{2}-1)\times d^{2}$ coefficient matrices, $F_{n}$, whose entries are $\{f_{x,y}^{n}\}$. We define $x$-th row and $y$-th column of $F_{n}$ respectively as $\mathbf{r}_{x}^{n}:=(f_{x,0}^{n},f_{x,1}^{n},\ldots,f_{x,d^{2}-1}^{n})^{\intercal}$ and $\mathbf{c}_{y}^{n}:=(f_{1,y}^{n},f_{2,y}^{n},\ldots,f_{d^{2}-1,y}^{n})^{\intercal}$ such that $F_{n}=[\mathbf{r}_{1}^{n},\mathbf{r}_{2}^{n},\ldots,\mathbf{r}_{d^{2}-1}^{n}]^{\intercal}=[\mathbf{c}_{0}^{n},\mathbf{c}_{1}^{n},\ldots,\mathbf{c}_{d^{2}-1}^{n}]$. In terms of $F_{n}$, real-valuedness condition is re-written as $F_{n}=F_{d-n}^{*}$.
Maximal value of $S$ under LHV theory is $L:=\max_{\mathbf{p}\in\mathcal{L}}S(\mathbf{p})$. Violation of the Bell inequality by quantum mechanics, $L<Q$, is shown with the maximal quantum value, $Q:=\max_{\mathbf{p}\in\mathcal{Q}}S(\mathbf{p})$. Quantum mechanical expectation can be evaluated with the operator,
$\mathcal{B}=\sum_{n}\sum_{x,y}f_{x,y}^{n}A_{x}^{n}\otimes B_{y}^{n}.$ (3)
Here, we consider the quantum measurements $A_{x}=\sum_{\alpha}\omega^{\alpha}G_{\alpha}^{x}$ and $B_{y}=\sum_{\beta}\omega^{\beta}H_{\beta}^{y}$ where $G_{\alpha}^{x}$, $H_{\beta}^{y}$ are rank-one projectors.
III Main result
We outline the method to derive the condition on the correlation matrix, $F_{n}$, such that maximal quantum value is obtained by generalized EBI correlation. SOS decomposition technique is employed for analytic derivation of maximal quantum value. Consider the shifted operator, ¯$\bar{\mathcal{B}}:=Q\mathds{I}-\mathcal{B}$ and the operators $P_{y}^{n}:=D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}$ defined with a real constant $Q\in\mathbb{R}$. The matrix $\mathds{I}:=\mathds{1}\otimes\mathds{1}$ is the $d^{2}\times d^{2}$ identity matrix, and $D_{y}^{n}:=\sum_{x}f_{x,y}^{n}A_{x}^{n}$. The SOS decomposition ¯$\bar{\mathcal{B}}=\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}\geq 0$ directly implies $\mathcal{B}\right>\leq Q$. Substituting $P_{y}^{n}$ into $\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}$ derives
$\mathcal{B}\leq\frac{1}{2}\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}\otimes\mathds{1}+\frac{1}{2}\sum_{n,y}\mathds{1}\otimes B_{y}^{n}(B_{y}^{n})^{\dagger}.$ (4)
In the derivation of (4), it is used the Hermiticity of $\mathcal{B}$, i.e. $\mathcal{B}=\sum_{n,y}D_{y}^{n}\otimes B_{y}^{n}=\sum_{n,y}(D_{y}^{n})^{\dagger}\otimes(B_{y}^{n})^{\dagger}$. The decomposition condition, ¯$\bar{\mathcal{B}}=\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}$, is obtained by deriving the RHS of (4) as $Q\mathds{I}$. Saturation condition is $\operatorname{tr}[\bar{\mathcal{B}}\rho]=0$ with any valid quantum realization. It implies that $Q$ is attainable upper bound. For all $y,n$, $P_{y}^{n}\rho=0$ is equivalent to the saturation condition [19]. Importantly, this condition is satisfied by the relation $D_{y}^{n}=(B_{y}^{n})^{*}$ on the support of the maximally entangled state as $P_{y}^{n}\left|\phi_{d}^{+}\right>=[D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}]\left|\phi_{d}^{+}\right>=0$. Lemma $M\otimes\mathds{1}\left|\phi_{d}^{+}\right>=\mathds{1}\otimes M^{\intercal}\left|\phi_{d}^{+}\right>$ is used in the second equation.
Our strategy is to solve the decomposition and saturation condition by varying $f_{x,y}^{n}$ with pre-assumed generalized EBI correlation. To this end, we define the optimal observables satisfying conditions (ii) and (iii),
$A_{dp+q}$ $=\tau W_{dp+q}\quad\,\forall p,q$ (5)
$B_{y}$ $=W_{y}B_{0}W_{y}^{\dagger}\quad\forall y$ (6)
where $\tau:=\omega^{pq\delta(d,2)/2}$ [39] defined with Kronecker delta $\delta$ and the parameterization $x=dp+q$ is used in Eq.(5). The index $x\in(0,d^{2})$ is now expressed with $p,q\in[0,d)$ satisfying $(p,q)\neq(0,0)$ and $y\in[0,d^{2})$. Alice’s observables in (5) are given as $d^{2}-1$ WH operators. Bob’s observables are generated from $B_{0}$ as Eq.(6). An important part of our method is to construct $B_{0}$ such that it has the fiducial state as its eigenstate. This choice of $B_{0}$ makes it possible to realize WH group covariant SICs on Bob’s eigenspace. Consider $B_{0}=\sum_{\beta}\omega^{\beta}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|$ where the eigenspace is defined as $\Phi(\varphi):=\{\left|\varphi_{\beta}\right>|\beta\in[0,d)\}$, a set of $d$ orthonormal states including the fiducial state $\left|\varphi\right>$. Without the loss of generality, one can let $\left|\varphi\right>:=\left|\varphi_{d-1}\right>$. Then we have,
$B_{y}$ $=\sum_{\beta}\omega^{\beta}W_{y}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|W_{y}^{\dagger}$ (7)
where the eigenbases of $B_{y}$ are now $\{W_{y}\left|\varphi_{\beta}\right>|\beta\in[0,d)\}$. Consequently, the set of $(d-1)$-th eigenstates of all the Bob’s observables comprises the $d^{2}$ WH group covariant SICs, $\{W_{y}\left|\varphi\right>|y\in[0,d^{2})\}$.
[…]
$\tilde{Q}:=\frac{1}{2}\left[\sum_{n}\lVert F_{n}\rVert^{2}+d^{2}(d-1)\right]$ (8)
when $\mathbf{r}_{a}^{n}\cdot\mathbf{r}_{x}^{n}=0$ for all cases of $a\neq x$.
Second, we solve the saturation condition with optimal realization, $\left|\phi_{d}^{+}\right>$ and observables (5) and (7). Saturation condition is given as $(B_{y}^{n})^{*}=D_{y}^{n}=\sum_{x}f_{x,y}^{n}W_{x}^{n}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ where $\mathbf{A}_{n}:=\tau^{n}(W_{1}^{n},W_{2}^{n},\ldots,W_{d^{2}-1}^{n})^{\intercal}$. The solution $\mathbf{c}_{y}^{n}$ completely derives the coefficient matrices we after. In sum, we have to solve $(B_{y}^{n})^{*}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ about $\{f_{x,y}^{n}\}$ under constraint $\mathbf{r}_{a}^{n}\cdot\mathbf{r}_{x}^{n}=0$ $\forall a\neq x$. The solution is derived in the below theorem.
Theorem. Consider prime $d$, for which WH group covariant SICs is defined. Then, there always exists the expression $S$ defined with,
$f_{dp+q,dr+s}^{n}=\omega^{-n(ps+qr)}f_{dp+q,0}^{n}$ (9)
for $p,q,r,s\in[0,d)$, $dp+q\neq 0$ that satisfies $Q=\tilde{Q}$ with generalized EBI correlation.
Proof.—Consider optimal $B_{0}=\sum_{\beta=0}^{d-1}\omega^{\beta}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|$ defined with eigenspace $\Phi(\varphi)$ containing $\left|\varphi\right>_{d-1}:=\left|\varphi\right>$. This choice of $B_{0}$ realizes SICs in the eigenspace as explained. Saturation condition is re-written with (6) as
$(B_{y}^{n})^{*}$ $=(W_{y}B_{0}^{n}W_{y}^{\dagger})^{*}$
$=\mathbf{c}_{0}^{n}\cdot(W_{y}^{*}\mathbf{A}_{n}W_{y}^{\intercal})$
$=\sum_{p,q}\omega^{-n(ps+qr)}f_{dp+q,0}^{n}(\mathbf{A}_{n})_{dp+q}$
$=\sum_{p,q}f_{dp+q,dr+s}^{n}(\mathbf{A}_{n})_{dp+q}$ (10)
where the notation $(\cdot)_{x}$ stands for the $x$-th element. The second equation is derived from saturation condition, $B_{0}^{n}=(\mathbf{c}_{0}^{n}\cdot\mathbf{A}_{n})^{*}$. Substituting $(\mathbf{A}_{n})_{dp+q}=(\tau X^{p}Z^{q})^{n}$ and $W_{y}=W_{dr+s}=X^{r}Z^{s}$ into RHS of the second equation and suitably applying the canonical commutation relation, $Z^{q}X^{p}=\omega^{pq}X^{p}Z^{q}$, derives the third equation. We used the parameterizations $x=dp+q$ and $y=dr+s$ for the substitution. The equation (10) is obtained from $(B_{dr+s}^{n})^{*}=\mathbf{c}_{dr+s}^{n}\cdot\mathbf{A}_{n}$ and it results in Eq.(9).
Coefficients in Eq.(9) also satisfies the decomposition condition. Substituting (9) into $\mathbf{r}_{a}^{n}\cdot\mathbf{r}_{x}^{n}=\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}$, one can find that it reduces to 0 for all $n$ if $a\neq x$. Finally, showing the saturation condition is always solvable completes the theorem. The elements of $\mathbf{A}_{n}$ is the unitary operator bases for traceless operator [34]. It means that the traceless operator $(B_{0}^{n})^{*}$ can always be expressed as $\mathbf{c}_{0}^{n}\cdot\mathbf{A}_{n}$ with suitable $\mathbf{c}_{0}^{n}$. $\blacksquare$
The above theorem states that the correlation function maximized by generalized EBI correlation always can be found from Eq.(9) with WH group covariant SICs. One has freedom in the choice of SICs, thus $B_{0}$, which determines the coefficients $\{f_{x,0}^{n}\}$ with $(B_{0}^{n})^{*}=\sum_{x}f_{x,0}^{n}(\mathbf{A}_{n})_{x}$. One can construct different correlation functions by varying $\{f_{x,0}^{n}\}$ of Eq.(9) with different choices of SICs. We also note that establishing $\{A_{x}\}$ as unitary operator bases [34] for traceless operator is advantageous for restricting the number of measurements, as the bases are minimum number of unitary operators that express, in linear combination form, an arbitrary traceless unitary operator, in our case $B_{0}^{n}$.
IV Derivation of Bell inequality
IV.1 Gisin’s elegant Bell inequality
[…]
IV.2 Bell inequality for two-qutrit system
We derive a Bell inequality maximally violated by generalized EBI correlation in $d=3$. For $d=3$, there are infinitely many fiducial states generating WH covariant SICs, for example $\left|\psi\right>=1/\sqrt{2}(0,1,-\omega^{\theta})$ for $20\leq\theta\leq 1/$ 2 [41]. We choose the one with $\theta=0$, $\left|\varphi\right>:=1/\sqrt{2}(0,1,-1)=1/\sqrt{2}\left(\left|1\right>-\left|2\right>\right)$ which is known to generate the Hesse SICs [42]. With the set of orthonormal bases $\left|\varphi_{0}\right>:=\left|0\right>$, $\left|\varphi_{1}\right>:=1/\sqrt{2}(\left|1\right>+\left|2\right>)$, and $\left|\varphi_{2}\right>:=\left|\varphi\right>$, optimal $B_{0}$ is
$B_{0}=\sum_{\beta}\omega^{\beta}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|=\begin{pmatrix}1&0&0\\ 0&-1/2&i\sqrt{3}/2\\ 0&i\sqrt{3}/2&-1/2\end{pmatrix}.$ (11)
It is worth mentioning here that our choice of $\left|\varphi_{1}\right>$ is also a fiducial state generating WH covariant SICs. By solving the saturation condition of SOS decomposition, $(B_{0})^{*}=\mathbf{c}_{0}^{1}\cdot\mathbf{A}_{1}$, the first column of $F_{1}$ is easily obtained as $\mathbf{c}_{0}^{1}=(f_{1,0}^{1},f_{2,0}^{1},\ldots,f_{8,0}^{1})^{\intercal}=(1/2)(1,1,\lambda,\nu,\mu,\lambda,\mu,\nu)^{\intercal}$ with $\lambda:=-i/\sqrt{3}$, $\mu:=\omega\lambda$ and $\nu:=\omega^{2}\lambda$. Substituting $\mathbf{c}_{0}$ into Eq.(9) and using the real-valuedness condition $F_{2}=F_{1}^{*}$, the coefficients are fully generated as in [40]. The derived coefficients can be further simplified in a functional form, $f_{dp+q,0}^{n}=1/2$ for $p=0$ and $f_{dp+q,0}^{n}=\omega^{-n(pq+3/2)+3/4}/(2\sqrt{3})$ for $p\neq 0$.
Substituting the derived coefficients into the expression (2), a Bell inequality maximally violated by generalized EBI correlation in two-qurtrit system is derived as,
$S_{3}:=\sum_{n=1}^{2}\sum_{\begin{subarray}{c}p,q,r,s=0\\ 3p+q\neq 0\end{subarray}}^{2}\omega^{-n(ps+qr)}f_{3p+q,0}^{n}\left\langle A_{3p+q}^{n}B_{3r+s}^{n}\right\rangle\leq 15.$ (12)
The upper bound $L=15$ can be obtained in the set of local deterministic correlations satisfying $P(\alpha\beta|xy)=1$ for certain combination of $\alpha,\beta$ and otherwise the probability is 0. The optimal strategy is found as, for all $x$, assigning $P(00|xy)$, $P(01|xy)$, $P(02|xy)$ as 1 respectively for $y=0,1,2$, $y=3,6$, and $y=4,5,7,8$. Violation is obtained from the maximal quantum value, $\tilde{Q}=18$, evaluated from (8) with $\lVert F_{n}\rVert=9$ for $n=1,2$. The upper bound is attained by two parties performing measurements, $A_{x}=W_{x}$ and $B_{y}=W_{y}B_{0}W_{y}^{\dagger}$, on the two-qutrit maximally entangled state when $x\in(0,9)$ and $y\in[0,9)$. As expected, $2d^{2}$ eigenstates, $W_{y}\left|\varphi_{1}\right>$ and $W_{y}\left|\varphi_{2}\right>$ for all $y$, comprise dual sets of $d^{2}$ SICs and eigenstates of four optimal observables for $x=1,3,4,5$, which are given as $Z,X,XZ,XZ^{2}$, defines the complete sets of MUBs [40].
We remark that the Bell inequality defined in (12) shows stronger violation than known Bell inequalities of similar classes. The critical visibility $\nu^{c}$, the smallest visibility $\nu$ of the isotropic state $\rho=\nu\left|\phi_{d}^{+}\right>\left<\phi_{d}^{+}\right|+(1-\nu)\mathds{I}/d^{2}$ to show the violation, is 83.33% for $S_{3}$. It is smaller than that of Bell inequalities including ones defined in $d=3$ for MUBs, 96.77% [19], 96.63% [20], for SICs, 96.41% [20], and original EBI, 86.6%. In addition, the number of settings required for Alice is 8 for $S_{3}$ and it is smaller than the case of known Bell inequality for $d=3$ SICs which requires $\binom{d^{2}}{2}=36$ settings for Alice when the number of Bob’s setting is same as 9 [20].
V Conclusion
We presented a method to construct correlation function for which the quantum maximum is always obtained in a simple analytic form from the generalized EBI correlation. Bell inequality for two-qutrit system derived in our method shows larger violation compared to known Bell inequalities of similar classes, including original EBI. Our work lays the foundation for the future studies for self-testing MUBs and SICs in high dimensional system and DI cryptographic protocols based on it.
Acknowledgements.
This research was supported by Korea Institute of Science and Technology Information(KISTI).(No. (KISTI)K24L4M1C2) and the National Research Council of Science & Technology(NST) grant by the Korea government (MSIT) (No. CAP22053-000). J. R. acknowledges support from the National Research Foundation (NRF) of Korea grant funded by the Korea Government (Grant No. NRF-2022M3K2A1083890).
Appendix A Symmetries on the coefficient matrices and optimal measurements
We present the coefficient matrix for $S_{3}$ and MUBs and SICs that gives the optimal violation. Although the symmetries in $S_{3}$ can be expressed with an analytic equation in the main text, the explicit form of the coefficient matrix, MUBs and SICs would show the symmetries in the forms that can be seen at a glance. The coefficient matrices are given as $8\times 9$ matrices satisfying real-valuedness condition, $F_{2}=F_{1}^{*}$ as follows
$F_{1}:=\frac{1}{2}$ $\begin{bmatrix}1&1&1&\omega^{2}&\omega^{2}&\omega^{2}&\omega&\omega&\omega\\ 1&1&1&\omega&\omega&\omega&\omega^{2}&\omega^{2}&\omega^{2}\\ \lambda&\nu&\mu&\lambda&\nu&\mu&\lambda&\nu&\mu\\ \nu&\mu&\lambda&\mu&\lambda&\nu&\lambda&\nu&\mu\\ \mu&\lambda&\nu&\nu&\mu&\lambda&\lambda&\nu&\mu\\ \lambda&\mu&\nu&\lambda&\mu&\nu&\lambda&\mu&\nu\\ \mu&\nu&\lambda&\lambda&\mu&\nu&\nu&\lambda&\mu\\ \nu&\lambda&\mu&\lambda&\mu&\nu&\mu&\nu&\lambda\end{bmatrix}$
when $\lambda:=-i/\sqrt{3}$, $\mu:=\omega\lambda$ and $\nu:=\omega^{2}\lambda$. One can find the first column generates the others following, $f_{dp+q,dr+s}^{1}=\omega^{-(ps+qr)}f_{3p+q,0}^{1}$.
[…]
$\mbox{SIC}_{1}$ $=\frac{1}{\sqrt{2}}\begin{bmatrix}0&1&1&0&\omega^{2}&\omega&0&\omega&\omega^{2}\\ 1&0&1&\omega&0&\omega^{2}&\omega^{2}&0&\omega\\ 1&1&0&\omega^{2}&\omega&0&\omega&\omega^{2}&0\end{bmatrix}\,$
$\mbox{SIC}_{2}$ $=\frac{1}{\sqrt{2}}\begin{bmatrix}0&-1&1&0&-\omega^{2}&\omega&0&-\omega&\omega^{2}\\ 1&0&-1&\omega&0&-\omega^{2}&\omega^{2}&0&-\omega\\ -1&1&0&-\omega^{2}&\omega&0&-\omega&\omega^{2}&0\end{bmatrix}$
where each column of $\mbox{SIC}_{1}$ and $\mbox{SIC}_{2}$ defines SICs generated from $\left|\varphi_{1}\right>$ and $\left|\varphi_{2}\right>$ respectively. On the other side, the eigenstates of Alice’s optimal observables $A_{1}=Z$, $A_{3}=X$, $A_{4}=XZ$, $A_{5}=XZ^{2}$ respectively define the complete mutually unbiased bases,
$\mbox{MUB}_{1}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix},\,\mbox{MUB}_{2}=\frac{1}{\sqrt{3}}\begin{bmatrix}1&\omega^{2}&\omega\\ 1&\omega&\omega^{2}\\ 1&1&1\end{bmatrix}$
$\mbox{MUB}_{3}=\frac{1}{\sqrt{3}}\begin{bmatrix}\omega^{2}&\omega&1\\ \omega^{2}&1&\omega\\ 1&1&1\end{bmatrix},\,\mbox{MUB}_{4}=\frac{1}{\sqrt{3}}\begin{bmatrix}\omega&1&\omega^{2}\\ \omega&\omega^{2}&1\\ 1&1&1\end{bmatrix}.$
We note that the eigenvectors of other settings $A_{2},A_{6},A_{7},A_{8}$ also define the above four MUBs. It can be easily shown from the fact that they are unitary and up to phase equivalent to $A_{1}^{2},A_{3}^{2},A_{5}^{2},A_{4}^{2}$ respectively.
Appendix B Derivation of the quantum bound
Detailed version of the derivation of quantum upper bound, $\tilde{Q}$, is to be presented here. The SOS decomposition $\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}\geq 0$ can be modified as
$\mathcal{B}\leq\frac{1}{2}\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}\otimes\mathds{1}+\frac{1}{2}\sum_{n,y}\mathds{1}\otimes B_{y}^{n}(B_{y}^{n})^{\dagger}$ (13)
by substituting $P_{y}^{n}:=D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}$. The second term of the RHS of (13) is reduced to $d^{2}(d-1)/2\mathds{1}$ from the unitarity of $B_{y}^{n}$. The first term of the RHS is derived as the desirable form by substituting $D_{y}^{n}:=\sum_{x}f_{x,y}^{n}A_{x}^{n}$ into $\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}$, as follows,
$\sum_{n,y}\left(\sum_{a}f_{a,y}^{n}A_{a}^{n}\right)^{\dagger}\sum_{x}f_{x,y}^{n}A_{x}^{n}$
$=\sum_{n,a,x}\left[\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}\right](A_{a}^{n})^{\dagger}A_{x}^{n}$
$=\sum_{n,x,y}|f_{x,y}^{n}|^{2}\mathds{1}+\sum_{n}\sum_{a\neq x}\left[\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}\right](A_{a}^{n})^{\dagger}A_{x}^{n}$
$=\sum_{n}\lVert F_{n}\rVert^{2}\mathds{1}+\sum_{a\neq x}(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}(A_{a}^{n})^{\dagger}A_{x}^{n}$ (14)
where the Frobeniuns norm $\lVert F_{n}\rVert:=(\sum_{x,y}|f_{x,y}^{n}|^{2})^{1/2}$. Consequently, the upper bound in (13) is reduced to
$\tilde{Q}:=\frac{1}{2}\left[\sum_{n}\lVert F_{n}\rVert^{2}+d^{2}(d-1)\right]$ (15)
under conditions $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ for all $n,a\neq x$.

### v2

@@
 $\tilde{Q}:=\frac{1}{2}\left[\sum_{n}\lVert F_{n}\rVert^{2}+d^{2}(d-1)\right]$ (8)
-when $\mathbf{r}_{a}^{n}\cdot\mathbf{r}_{x}^{n}=0$ for all cases of $a\neq x$.
+when $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ for all cases of $a\neq x$.
 Second, we solve the saturation condition with optimal realization, $\left|\phi_{d}^{+}\right>$ and observables (5) and (7).
@@
 The solution $\mathbf{c}_{y}^{n}$ completely derives the coefficient matrices we after.
-In sum, we have to solve $(B_{y}^{n})^{*}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ about $\{f_{x,y}^{n}\}$ under constraint $\mathbf{r}_{a}^{n}\cdot\mathbf{r}_{x}^{n}=0$ $\forall a\neq x$.
+In sum, we have to solve $(B_{y}^{n})^{*}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ about $\{f_{x,y}^{n}\}$ under constraint $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ $\forall a\neq x$.
 The solution is derived in the below theorem.
@@
 Coefficients in Eq.(9) also satisfies the decomposition condition.
-Substituting (9) into $\mathbf{r}_{a}^{n}\cdot\mathbf{r}_{x}^{n}=\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}$, one can find that it reduces to 0 for all $n$ if $a\neq x$.
+Substituting (9) into $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}$, one can find that it reduces to 0 for all $n$ if $a\neq x$.
 Finally, showing the saturation condition is always solvable completes the theorem.
@@
 The derived coefficients can be further simplified in a functional form, $f_{dp+q,0}^{n}=1/2$ for $p=0$ and $f_{dp+q,0}^{n}=\omega^{-n(pq+3/2)+3/4}/(2\sqrt{3})$ for $p\neq 0$.
-Substituting the derived coefficients into the expression (2), a Bell inequality maximally violated by generalized EBI correlation in two-qurtrit system is derived as,
+Substituting the derived coefficients into (2), a Bell inequality maximally violated by generalized EBI correlation in two-qurtrit system is derived as,
 $S_{3}:=\sum_{n=1}^{2}\sum_{\begin{subarray}{c}p,q,r,s=0\\ 3p+q\neq 0\end{subarray}}^{2}\omega^{-n(ps+qr)}f_{3p+q,0}^{n}\left\langle A_{3p+q}^{n}B_{3r+s}^{n}\right\rangle\leq 15.$ (12)
@@
 Bell inequality for two-qutrit system derived in our method shows larger violation compared to known Bell inequalities of similar classes, including original EBI.
-Our work lays the foundation for the future studies for self-testing MUBs and SICs in high dimensional system and DI cryptographic protocols based on it.
+Our work lays the foundation for the future studies for self-testing MUBs and SICs in high dimensional system and DI quantum cryptographic protocols based on it.
 Acknowledgements.
@@
 Appendix B Derivation of the quantum bound
-Detailed version of the derivation of quantum upper bound, $\tilde{Q}$, is to be presented here. The SOS decomposition $\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}\geq 0$ can be modified as
+Detailed version of the derivation of quantum upper bound, $\tilde{Q}$, is presented in this section.
+As shown in the main text, the SOS decomposition $\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}\geq 0$ can be modified as
 $\mathcal{B}\leq\frac{1}{2}\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}\otimes\mathds{1}+\frac{1}{2}\sum_{n,y}\mathds{1}\otimes B_{y}^{n}(B_{y}^{n})^{\dagger}$ (13)

### v3

@@
-Designing Elegant Bell Inequalities
+Designing generalized elegant Bell inequalities in high dimension from a quantum bound
 Abstract
-Elegant Bell inequality is well known for its much exploited property, being maximally violated by maximal entanglement, mutually unbiased bases, and symmetric informationally complete positive operator-valued measure elements.
-It is the only one with such property known so far.
-We present a method to construct Bell inequalities with violation feature analogous to original elegant Bell inequality in high dimension from a simple analytic quantum bound.
-A Bell inequality with such feature is derived in three dimension for the first time.
-It shows larger violation than existing Bell inequalities of similar classes while requiring arguably small number of measurements.
-I Introduction
+Elegant Bell inequality is well known for its distinctive property, being maximally violated by maximal entanglement, mutually unbiased bases, and symmetric informationally complete positive operator-valued measure elements.
+Despite its significance in quantum information theory demonstrated based on its unique violation feature, it remains the only known one with the characteristic.
+We present a method to construct Bell inequalities with violation feature analogous to elegant Bell inequality in higher local dimension from a simple analytic quantum bound.
+A Bell inequality with the generalized violation feature is derived in three dimension for the first time.
+It exhibits larger violation than existing Bell inequalities of similar classes, including the original elegant Bell inequality, while requiring arguably small number of measurements.
+1 Introduction
 Performing local measurements on the shared system, distant subsystems can reveal a quantum correlation that cannot be reproduced by any local hidden variable theoretical explanation.
 Such nonlocality of the correlation can be tested via the violation of Bell inequalities [1, 2].
-What Bell inequality can tell us about the system goes beyond the scope of the mere existence of nonlocal correlation.
-It has been investigated the method to exploit Bell’s theorem for further certifying the properties of given system in so-called device-independent (DI) approach [3, 4].
+The implications of Bell inequalities extend beyond merely confirming the existence of nonlocal correlations.
+Researchers have investigated methods to exploit Bell’s theorem for certifying the properties of given system through the device-independent (DI) approach [3, 4, 5].
 This approach leads to various applications, including DI quantum key distribution [4, 6, 7, 8, 9] and DI certification of randomness [10, 11, 12].
-Reflecting on the aforementioned possibilities provided by Bell inequalities, it is of prime importance to derive them such that their maximal violations are obtained from desirable features, for example maximal entanglement [12, 13, 14, 15, 16, 17, 18, 19, 20], mutually unbiased bases (MUBs) [19, 20], or symmetric informationally complete POVM elements (SICs) [21, 20].
+Reflecting on the aforementioned possibilities provided by Bell inequalities, it is of prime importance to derive them such that their maximal violations are obtained from desirable features, for example maximal entanglement [13, 14, 15, 16, 17, 18, 19, 20, 21], mutually unbiased bases (MUBs) [20, 21], or symmetric informationally complete quantum measurements (SICs) [22, 21].
 As is well known, Clauser-Horne-Shimony-Holt Bell inequality is the case in which the maximal violation can be obtained from maximal entanglement and Pauli operators having MUBs as their eigenstates.
@@
 What makes EBI more interesting is that its violation is obtained from the case where the local measurements of each subsystem is constructed by MUBs and SICs respectively.
-As first remarked by Gisin, the geometrical symmetries in the optimal measurement settings of EBI are clearly revealed when they are mapped on Bloch’s sphere [21].
-Three eigenstates defining three projective measurements of Alice form three mutually orthogonal vectors on Bloch sphere and four eigenstates of Bob’s side make a tetrahedron [21].
-Original EBI has been used in several fields of quantum information, including DI randomness certification [22, 23], DI certification of non-projective measurement [24, 25], and for other tasks [26, 27, 28].
-Self-testing property of EBI has also been studied in [29].
-MUBs and SICs are important resources in quantum cryptography [30, 31, 32, 22, 23].
-There has been efforts to derive Bell inequalities for MUBs or SICs exploiting the advantage of high dimension [19, 20].
+As first remarked by Gisin, the geometrical symmetries in the optimal measurements of EBI are clearly revealed when they are mapped on Bloch’s sphere [22].
+Three eigenstates defining three projective measurements of Alice form three mutually orthogonal vectors on Bloch sphere and Bob’s four eigenstates make a tetrahedron [22].
+EBI has been applied for several tasks in quantum information field, including DI randomness certification [23, 24], DI certification of non-projective measurement [25, 26], and other tasks [27, 28, 29].
+Self-testing property of EBI is studied in [30].
+MUBs and SICs are not only relevant for the formulation of quantum mechanics [31], but also for the extensive quantum information processing tasks [23, 24, 32, 33, 34, 35, 36, 37, 38, 39, 40].
+There has been efforts to derive Bell inequalities for MUBs or SICs exploiting the high dimensional system [20, 21].
 For DI randomness certification, for which high dimension provides natural advantage [23], Bell inequalities for SICs has been used [21, 23, 24].
-EBI is distinguished from ones either for MUBs or SICs as it is optimally violated by both of them.
-We address the problem of deriving Bell inequalities with violation feature analogous to original EBI in high dimensions.
-The problem is closely related to the question of what benefit for Bell nonlocality would come from simultaneous consideration of the symmetries of MUBs and SICs.
-It is worth noting that although EBI is originally derived in the different context [33, 21], we focus on the generalization of the property, being maximally violated by maximal entanglement, MUBs and SICs.
-Such generalization of EBI has not much been investigated so far to our knowledge.
-We approach the raised problem by defining a generalized type of correlation as follows.
-Definition.
-For prime $d\geq 3$, generalized EBI correlation is defined as the bipartite correlation obtained from the quantum realization: (i) maximally entangled state $\left|\phi_{d}^{+}\right>:=(1/\sqrt{d})\sum_{\alpha=0}^{d-1}\left|\alpha\alpha\right>$, (ii) Alice’s $d^{2}-1$ observables comprising $\mathcal{W}_{d}\setminus\mathds{1}$ and (iii) Bob’s $d^{2}$ observables having $d^{2}$ SICs in their eigenspaces.
-In condition (ii), $\mathcal{W}_{d}:=\{X^{m}Z^{n}|m,n\in[0,d)\}$ is the set of $d^{2}$ Weyl-Heisenberg (WH) operators defined with $X:=\sum_{\alpha=0}^{d-1}\left|\alpha+1\right>\left<\alpha\right|$, $Z:=\sum_{\alpha=0}^{d-1}\omega^{\alpha}\left|\alpha\right>\left<\alpha\right|$ and $\omega:=e^{2\pi i/d}$.
+EBI is distinguished from those for either MUBs or SICs as it is optimally violated by both.
+Randomness certification is a notable application of EBI.
+DI certification scheme of optimal randomness from one entanglement bit is suggested based on EBI [23, 24].
+Considering the above, investigating the desirable feature of Bell inequalities analogous to EBI in high dimension is expected to be intriguing from the application perspective beyond its implications just for extending the mathematical symmetries.
+We address the problem of deriving Bell inequalities in higher dimensions with violation features analogous to EBI.
+This investigation is closely related to understanding how the simultaneous consideration of MUB and SIC symmetries could enhance Bell nonlocality.
+Although EBI originated in a different context [22, 41], our focus lies in generalizing its distinctive property of achieving maximal violation through maximal entanglement, MUBs, and SICs.
+Such generalization of EBI remains largely unexplored in the literature.
+This article is structured as follows.
+In Sec.
+2, we briefly introduce MUBs and SICs, key concepts of our result.
+Bell scenario and notations defining it are given in the same section.
+In Sec.
+3, the generalized violation feature inspired from EBI is defined.
+It is denoted under a term, generalized EBI correlation.
+In Sec.
+4, we present the main result deriving the form of correlation function maximized with the generalized EBI correlation.
+The corresponding quantum bound is derived in an analytic form.
+In Sec.
+5, to show the applicability of our result, we derive Gisin’s EBI and a novel Bell inequality for three dimensional generalized EBI correlation.
+We evaluate the violation of the Bell inequality in comparison with established results.
+2 Preliminaries
+2.1 MUBs and SICs
+We briefly introduce the definition of MUBs and SICs, along with their properties highly relevant to this work.
+Comprehensive reviews of MUBs and SICs, including their fundamental concepts and properties, can be found in [35, 42].
+2.1.1 MUBs
+Normalized bases $\{\left|a_{i}\right>\}_{i=0}^{d-1}$ and $\{\left|b_{j}\right>\}_{j=0}^{d-1}$ are said to be mutually unbiased when they satify
+|$a_{i}|b_{j}\right>\right|^{2}=\frac{1}{d}$ (1)
+for all combinations of $i$ and $j$.
+Problem of finding MUBs in all Hilbert space dimesion is still open.
+However, it is known that whenever the prime power dimension $d$ is considered there are at most $d+1$ MUBs [40], referred to as the complete set.
+The complete set of MUBs in prime $d$ dimension can be constructed using Weyl-Heisenberg (WH) group elements.
+The shift $X:=\sum_{\alpha=0}^{d-1}\left|\alpha+1\right>\left<\alpha\right|$ and the phase $Z:=\sum_{\alpha=0}^{d-1}\omega^{\alpha}\left|\alpha\right>\left<\alpha\right|$ operators generate the WH group $\{\omega^{n}X^{p}Z^{q}|n,p,q\in[0,d)\}$ with $\omega:=e^{2\pi i/d}$.
 The integer domain $\{0,1,\ldots,d-1\}$ is denoted by $[0,d)$ and similar abbreviation is to be used throughout this work.
-Note that $X^{0}Z^{0}=\mathds{1}$ is not included in the condition (ii).
-Conditions (i) and (iii) are natural generalization of optimal quantum realization of original EBI.
-As for the generalization of Alice’s optimal observables, two interpretations are possible.
-The set of Alice’s optimal observables for EBI, $\sigma_{x},\sigma_{y},\sigma_{z}$, [22] can be viewed either as “$d+1$ observables” defining $d+1$ sets of MUBs, or as “$d^{2}-1$ unitary operator bases” [34] for a traceless operator $\vec{a}\cdot\vec{\sigma}$ defined with a three-dimensional unit vector $\vec{a}$ and $\vec{\sigma}:=(\sigma_{x},\sigma_{y},\sigma_{z})$ when $d=2$.
-The above duality arises as $d+1=d^{2}-1$ when $d=2$.
-Condition (ii) clearly generalizes the latter.
-Under condition (ii), complete MUBs are realized in eigenspace of $d+1$ among $d^{2}-1$ observables of Alice, namely $Z$, $X$, $XZ$, …, $XZ^{d-1}$ [35].
-Moreover, the eigenbases of all $d^{2}-1$ observables of Alice comprises the $d-1$ same sets of complete $d+1$ MUBs.
-It can be easily seen from the fact that $XZ^{n}$ and $X^{m}Z^{mn}$ has the same eigenbases for all $m$ and $\mathcal{W}=\{X^{m}Z^{mn}|m,n\in[0,d)\}$ for prime $d$.
-In the following, we will further explain the reason for the choice of optimal observables of Alice as unitary operator bases.
-Figure 1: On the shared system, two subsystems, Alice and Bob, respectively performs one of $d^{2}-1$ and $d^{2}$ measurements outputting one of $d$ outcomes.
-Another insight that underlies our approach is to exploit $systematic$ and $versatile$ features of WH group covariant SICs.
-For brevity, consider the parameterization $W_{dp+q}:=X^{p}Z^{q}$, such that $\mathcal{W}_{d}=\{W_{j}|j\in[0,d^{2})\}$.
-Group covariant SICs is derived by acting group elements on a single state, so-called fiducial vector [36].
-For the case of WH group covariant SICs, group elements $W_{j}$ generates $d^{2}$ SICs on the support of fiducial state $\left|\varphi\right>$, $\{W_{j}\left|\varphi\right>|j\in[0,d^{2})\}$.
-Weyl-Heisenberg group covariance plays crucial role for satisfying the condition (iii) and provides systematic feature for our formalism.
-Versatility of our method is from the fact that it can be applied only if WH group covariant SICs for given local dimension exists.
-The existence of SICs are known in all dimension up to $d=193$ [37].
-More importantly, all group covariant SICs is covariant to WH group in prime dimensions [38].
-II Preliminaries
-We consider the measurement scenario of two subsystems Alice and Bob respectively performing one of possible $d^{2}-1$ and $d^{2}$ number of $d$ outcome measurements $A_{x}$ and $B_{y}$ on the shared system.
+Note that $X^{d}=Z^{d}=\mathds{1}$ and the canonical commutation relation (CCR) holds, i.e.
+$Z^{q}X^{p}=\omega^{pq}X^{p}Z^{q}$.
+A subset of WH group without phase factor also forms a group,
+$\mathcal{W}(d)=\{X^{m}Z^{n}|m,n\in[0,d)\}.$ (2)
+The complete MUBs can be obtained as the eigenvectors of $d+1$ elements in $\mathcal{W}(d)$, namely $Z$, $X$, $XZ$, …, $XZ^{d-1}$ [43].
+2.1.2 SICs
+Suppose a set of $d^{2}$ normalized states $\{\left|e_{i}\right>\}_{i=0}^{d^{2}-1}$ satisfying the relation
+|$e_{i}|e_{j}\right>\right|^{2}=\frac{1}{d+1}$ (3)
+for all $i\neq j$.
+The above constant overlap relations reflects the geometrical symmetry among $d^{2}$ states.
+They are equiangular lines in $\mathds{C}^{d}$ [42].
+The states define a symmetric informationally complete positive operator-valued measure (POVM), $\left\{\frac{1}{d}\left|e_{i}\right>\left<e_{i}\right||i\in[0,d^{2})\right\}$ with additional condition that all the POVM elements must be summed up to become identity operator.
+The POVM is informationally complete in the sense that its statistics characterizes the quantum state $\rho$ in complex $d$ dimension upon which it is carried on.
+The $d^{2}$ number of elements are necessary for a informationally complete POVM.
+In this article, a SICs specifically denotes the $d^{2}$ unit vectors $\{\left|e_{i}\right>\}_{i=0}^{d^{2}-1}$ satisfying the relation (3).
+It is interesting that SICs are also closely related to WH group.
+WH operator bases hence provide a suitable common ground for the analysis of MUBs and SICs in high dimension.
+It is the very reason why we adopt the WH bases.
+SIC-POVM is group covariant if there is a group $\{U_{i}\}$ having unitary operators as its elements and SICs is generated from operating the elements upon a single vector (or state), so-called fiducial vector $\left|f\right>$, i.e.
+when $\{U_{i}\left|f\right>\}$ is SICs [45].
+The group covariance of SICs plays a significant role in the problem of SICs generation [42].
+2.2 Bell Scenario
+We consider the measurement scenario of two subsystems Alice (Bob) performing one of possible $d^{2}-1$ ($d^{2}$) number of $d$ outcome measurements $A_{x}$ ($B_{y}$) on their shared system.
 Measurements are labelled with $x\in\{1,2,\ldots,d^{2}-1\}$ and $y\in\{0,1,\ldots,d^{2}-1\}$.
-Outcomes of $A_{x}$ and $B_{y}$ are respectively distinguished with indices $\alpha,\beta\in\{0,1,\ldots,d-1\}$.
+Outcomes of $A_{x}$ and $B_{y}$ are respectively distinguished by indices $\alpha\in\{0,1,\ldots,d-1\}$ and $\beta$ in the same domain.
 The scenario is illustrated in Fig.
@@
 Probability of obtaining the outcomes $\alpha$ and $\beta$ respectively from settings $x$ and $y$ is denoted by $p(\alpha\beta|xy)$.
-A statistics of the Bell test can be described with correlation, $\mathbf{p}:=\{P(\alpha\beta|xy)|\forall\alpha,\beta,x,y\}$.
+A statistics of the Bell test can be described with $\mathbf{p}:=\{p(\alpha\beta|xy)\}_{\alpha,\beta,x,y}$, a correlation.
 Quantum correlation allows
 $p(\alpha\beta|xy)=\operatorname{tr}\left[\rho_{AB}(G_{\alpha}^{x}\otimes H_{\beta}^{y})\right]$
-where $\rho_{AB}$ is the bipartite density matrix of Alice and Bob.
-Measurement operators, $\{G_{\alpha}^{x}\}$ and $\{H_{\beta}^{y}\}$, define the measurements of Alice and Bob respectively.
-Bell inequality is used to show that there is quantum correlation which does not allow the local hidden variable (LHV) theoretical description of the correlation [1].
-Geometrically, LHV correlation region forms a polytope, $\mathcal{L}$, and quantum region, $\mathcal{Q}$, is known to include it [2].
-The generic form of so-called Bell expression [2] without marginal probabilities can be defined as,
+where $\rho_{AB}$ is a bipartite density matrix of Alice and Bob.
+The operators, $\{G_{\alpha}^{x}\}$ and $\{H_{\beta}^{y}\}$, define the measurements of Alice and Bob respectively.
+Bell inequality is used to show that there is quantum correlation, which does not allow the local hidden variable (LHV) theoretical description of the correlation [1].
+Geometrically, LHV correlation region forms a polytope, $\mathcal{L}$, and quantum region, $\mathcal{Q}$, includes it [2].
+The generic form of so-called Bell expression [2] without marginal terms can be defined as
 $S$ $=\sum_{\alpha,\beta}\sum_{x,y}g_{x,y}^{\alpha,\beta}P(\alpha\beta|xy)$ (4)
 $=\sum_{n}\sum_{x,y}f_{x,y}^{n}\langle A_{x}^{n}B_{y}^{n}\rangle.$ (5)
-where $n\in(1,d)$ and $g_{x,y}^{\alpha,\beta}\in\mathbb{R}$ is the real weight for each probability.
+where $n\in[1,d)$ and $g_{x,y}^{\alpha,\beta}\in\mathbb{R}$ is the weight for each probability.
 For simplicity, the summations over all possible values are to be denoted without limits.
 The correlator $\langle A_{x}^{n}B_{y}^{n}\rangle=\sum_{\alpha,\beta=0}^{d-1}\omega^{n(\alpha+\beta)}P(\alpha\beta|xy)$, where the outcomes of the measurements $A_{x}$ and $B_{y}$ are set as $\omega^{\alpha},\omega^{\beta}\in\{1,\omega,\omega^{2},\ldots,\omega^{d-1}\}$.
-And the coefficient $g_{x,y}^{\alpha,\beta}:=\sum_{n}f_{x,y}^{n}\omega^{n(\alpha+\beta)}$ The condition $f_{x,y}^{d-n}:=(f_{x,y}^{n})^{*}$ guarantees the real-valuedness of $S$.
-A expression is specified by a $(d^{2}-1)\times d^{2}$ coefficient matrices, $F_{n}$, whose entries are $\{f_{x,y}^{n}\}$.
-We define $x$-th row and $y$-th column of $F_{n}$ respectively as $\mathbf{r}_{x}^{n}:=(f_{x,0}^{n},f_{x,1}^{n},\ldots,f_{x,d^{2}-1}^{n})^{\intercal}$ and $\mathbf{c}_{y}^{n}:=(f_{1,y}^{n},f_{2,y}^{n},\ldots,f_{d^{2}-1,y}^{n})^{\intercal}$ such that $F_{n}=[\mathbf{r}_{1}^{n},\mathbf{r}_{2}^{n},\ldots,\mathbf{r}_{d^{2}-1}^{n}]^{\intercal}=[\mathbf{c}_{0}^{n},\mathbf{c}_{1}^{n},\ldots,\mathbf{c}_{d^{2}-1}^{n}]$.
-In terms of $F_{n}$, real-valuedness condition is re-written as $F_{n}=F_{d-n}^{*}$.
+Equivalence of the expressions (4) and (5) are given by the relation $g_{x,y}^{\alpha,\beta}:=\sum_{n}f_{x,y}^{n}\omega^{n(\alpha+\beta)}$ The condition $f_{x,y}^{d-n}:=(f_{x,y}^{n})^{*}$ guarantees the real-valuedness of the expression.
+A possible expression is specified by a $(d^{2}-1)\times d^{2}$ coefficient matrices, $F_{n}$, whose entry $(F_{n})_{x,y}$ is defined as $f_{x,y}^{n}$.
+We define $x$-th row and $y$-th column of $F_{n}$ respectively with column vectors $\mathbf{r}_{x}^{n}:=(f_{x,0}^{n},f_{x,1}^{n},\ldots,f_{x,d^{2}-1}^{n})^{\intercal}$ and $\mathbf{c}_{y}^{n}:=(f_{1,y}^{n},f_{2,y}^{n},\ldots,f_{d^{2}-1,y}^{n})^{\intercal}$ such that $F_{n}=[\mathbf{r}_{1}^{n},\mathbf{r}_{2}^{n},\ldots,\mathbf{r}_{d^{2}-1}^{n}]^{\intercal}=[\mathbf{c}_{0}^{n},\mathbf{c}_{1}^{n},\ldots,\mathbf{c}_{d^{2}-1}^{n}]$.
+The real-valuedness condition can be re-written as $F_{n}=F_{d-n}^{*}$, in terms of coefficient matrices.
 Maximal value of $S$ under LHV theory is $L:=\max_{\mathbf{p}\in\mathcal{L}}S(\mathbf{p})$.
@@
 Here, we consider the quantum measurements $A_{x}=\sum_{\alpha}\omega^{\alpha}G_{\alpha}^{x}$ and $B_{y}=\sum_{\beta}\omega^{\beta}H_{\beta}^{y}$ where $G_{\alpha}^{x}$, $H_{\beta}^{y}$ are rank-one projectors.
-III Main result
-We outline the method to derive the condition on the correlation matrix, $F_{n}$, such that maximal quantum value is obtained by generalized EBI correlation.
-SOS decomposition technique is employed for analytic derivation of maximal quantum value.
-Consider the shifted operator, ¯$\bar{\mathcal{B}}:=Q\mathds{I}-\mathcal{B}$ and the operators $P_{y}^{n}:=D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}$ defined with a real constant $Q\in\mathbb{R}$.
-The matrix $\mathds{I}:=\mathds{1}\otimes\mathds{1}$ is the $d^{2}\times d^{2}$ identity matrix, and $D_{y}^{n}:=\sum_{x}f_{x,y}^{n}A_{x}^{n}$.
+[…]
+[scale=0.13]fig_scenario.jpg
+Figure 1: The Bell scenario considers the $d^{2}-1$ and $d^{2}$ inputs in local systems outputting one of $d$ outcomes.
+3 Generalized EBI correlation
+In the introduction, the problem of deriving Bell inequalities analogous to EBI in high dimension is raised. We approach the problem by defining an optimal correlation as follows.
+Definition. For prime $d\geq 3$, a generalized EBI correlation is defined as the bipartite correlation obtained from the quantum realization:
+{adjustwidth}
+.5em0pt (i) $\left|\phi_{d}^{+}\right>:=(1/\sqrt{d})\sum_{\alpha=0}^{d-1}\left|\alpha\alpha\right>$
+(ii) Alice’s $d^{2}-1$ observables comprising $\mathcal{W}_{d}\setminus\mathds{1}$
+(iii) Bob’s $d^{2}$ observables respectively having each one of the $d^{2}$ unit vectors satisfying (3) as an eigenstate.
+Conditions (i)-(iii) define a possible generalization of optimal quantum realization of EBI.
+Condition (i) is a two-qudit maximally entangled state.
+Condition (iii) guarantees the realization of at least one SICs in the Bob’s eigenspace.
+These conditions are straightforwardly derived from the optimal realization of EBI.
+As for the generalization of Alice’s optimal observables, two interpretations are possible.
+The set of Alice’s optimal observables for EBI, $\sigma_{x},\sigma_{y},\sigma_{z}$, [23] can be viewed either as “$d+1$ observables” defining $d+1$ sets of MUBs or as “$d^{2}-1$ unitary operator bases” [31].
+Condition (ii) clearly generalizes the latter.
+Note that a traceless operator can be expressed as a familiar form $\vec{a}\cdot\vec{\sigma}$ with the unitary operator bases, $\vec{\sigma}:=(\sigma_{x},\sigma_{y},\sigma_{z})$, when $d=2$ and $\vec{a}$ is a three dimensional unit vector.
+The above duality arises as $d+1=d^{2}-1$ when $d=2$.
+Excluding trivial case $X^{0}Z^{0}=\mathds{1}$ makes the number of observables $d^{2}-1$ rather than $d^{2}$ in condition (ii).
+Complete MUBs are realized in eigenspace of $d+1$ among $d^{2}-1$ observables of Alice, namely $Z$, $X$, $XZ$, …, $XZ^{d-1}$ [43], under condition (ii).
+Moreover, the eigenstates of all $d^{2}-1$ observables of Alice falls into one of eigenstates of $Z$, $X$, $XZ$, …, $XZ^{d-1}$ under condition (ii).
+In other words, the eigenspaces of $d^{2}-1$ observables are $d-1$ folded eigenspaces of $d+1$ complete MUBs.
+(See appendix A for the detailed discussion.)
+WH operators play a pivotal role in not only deriving Alice’s optimal observable but also for Bob’s ones.
+The explicit derivation is given in the following section.
+Here, we briefly convey the advantages brought by exploiting WH group covariant SICs for satisfying condition (iii).
+For brevity, consider the parameterization $W_{dp+q}:=X^{p}Z^{q}$ such that $\mathcal{W}(d)=\{W_{j}|j\in[0,d^{2})\}$.
+For example, $W_{5}=XZ^{2}$ when $d=3$.
+With the parameterization, WH group covariant SICs are $\{W_{j}\left|\varphi\right>|j\in[0,d^{2})\}$ for a fiducial state $\left|\varphi\right>$.
+Group covariant SICs enable one to satisfy the condition (iii) in a systematic way.
+The condition is satisfied by assigning each one of $d^{2}$ states, $W_{j}\left|\varphi\right>$, to a different eigenstate of Bob’s $j$-th measurement.
+Consequently, Bob’s optimal measurements defined in condition (iii) are derived neatly as (9), in the following section.
+The WH group covariant SICs also provide versatile nature for our framework.
+It is based on the fact that our method for deriving the quantum bound can always be applied if WH group covariant SICs of given dimension exists.
+The existence of SICs is known in all dimension up to $d=193$ [44].
+More importantly, all group covariant SICs are covariant to WH group in prime dimension [46] that we mainly consider.
+In light of the above discussion, exploiting WH group covariance of SICs is arguably reasonable to derive and analyze the Bell nonlocality for generalized EBI correlation.
+In the following section, we elaborate on how the WH operator bases also offer advantages in the optimization of the quantum correlation.
+4 Main result
+We present a method to construct Bell expressions, always maximized by generalized EBI correlation if only the WH group covariant SICs is defined in a given dimension.
+Sum-of-squares (SOS) decomposition technique is used for deriving maximal quantum value in an analytic form.
+Consider the shifted operator, ¯$\bar{\mathcal{B}}:=Q\mathds{I}-\mathcal{B}$ and operators $P_{y}^{n}:=D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}$ where $Q\in\mathbb{R}$ is a real constant, $D_{y}^{n}:=\sum_{x}f_{x,y}^{n}A_{x}^{n}$, and $\mathds{1}$ is a $d\times d$ identity matrix.
+The identity matrix acting on the composite system is denoted by $\mathds{I}:=\mathds{1}\otimes\mathds{1}$.
 The SOS decomposition ¯$\bar{\mathcal{B}}=\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}\geq 0$ directly implies $\mathcal{B}\right>\leq Q$.
-Substituting $P_{y}^{n}$ into $\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}$ derives
-$\mathcal{B}\leq\frac{1}{2}\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}\otimes\mathds{1}+\frac{1}{2}\sum_{n,y}\mathds{1}\otimes B_{y}^{n}(B_{y}^{n})^{\dagger}.$ (4)
-In the derivation of (4), it is used the Hermiticity of $\mathcal{B}$, i.e.
-$\mathcal{B}=\sum_{n,y}D_{y}^{n}\otimes B_{y}^{n}=\sum_{n,y}(D_{y}^{n})^{\dagger}\otimes(B_{y}^{n})^{\dagger}$.
-The decomposition condition, ¯$\bar{\mathcal{B}}=\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}$, is obtained by deriving the RHS of (4) as $Q\mathds{I}$.
-Saturation condition is $\operatorname{tr}[\bar{\mathcal{B}}\rho]=0$ with any valid quantum realization.
-It implies that $Q$ is attainable upper bound.
-For all $y,n$, $P_{y}^{n}\rho=0$ is equivalent to the saturation condition [19].
-Importantly, this condition is satisfied by the relation $D_{y}^{n}=(B_{y}^{n})^{*}$ on the support of the maximally entangled state as $P_{y}^{n}\left|\phi_{d}^{+}\right>=[D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}]\left|\phi_{d}^{+}\right>=0$.
-Lemma $M\otimes\mathds{1}\left|\phi_{d}^{+}\right>=\mathds{1}\otimes M^{\intercal}\left|\phi_{d}^{+}\right>$ is used in the second equation.
-Our strategy is to solve the decomposition and saturation condition by varying $f_{x,y}^{n}$ with pre-assumed generalized EBI correlation.
-To this end, we define the optimal observables satisfying conditions (ii) and (iii),
+Substituting $P_{y}^{n}$ into the decomposition derives
+$\frac{1}{2}\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}\otimes\mathds{1}+\frac{1}{2}\sum_{n,y}\mathds{1}\otimes B_{y}^{n}(B_{y}^{n})^{\dagger}=Q\mathds{I}.$ (7)
+In the derivation, the Hermiticity of $\mathcal{B}$, i.e.
+$\mathcal{B}=\sum_{n,y}D_{y}^{n}\otimes B_{y}^{n}=\sum_{n,y}(D_{y}^{n})^{\dagger}\otimes(B_{y}^{n})^{\dagger}$, is used.
+Hereafter, the Eq.(7) is to be referred to as ‘decomposition condition’.
+We define ‘saturation condition’ as $\operatorname{tr}[\bar{\mathcal{B}}\rho]=0$ for any valid quantum realization.
+It implies that $Q$ is attainable from quantum mechanics.
+It is known that $P_{y}^{n}\rho=0\forall y,n$ is equivalent to the saturation condition [20].
+Importantly, this condition is satisfied by $D_{y}^{n}=(B_{y}^{n})^{*}$ on a maximally entangled state as $P_{y}^{n}\left|\phi_{d}^{+}\right>=[D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}]\left|\phi_{d}^{+}\right>=[D_{y}^{n}-(B_{y}^{n})^{*}]\otimes\mathds{1}\left|\phi_{d}^{+}\right>$.
+An identity, $M\otimes\mathds{1}\left|\phi_{d}^{+}\right>=\mathds{1}\otimes M^{\intercal}\left|\phi_{d}^{+}\right>\forall M$, is used in the second equation.
+This condition reveals the relation between optimal measurements on both subsystems.
+Our strategy is to solve the decomposition and saturation conditions simultaneously by varying the coefficients of Bell expression $f_{x,y}^{n}$ with a given generalized EBI correlation.
+To this end, we define the optimal measurements obeying conditions (ii) and (iii),
 $A_{dp+q}$ $=\tau W_{dp+q}\quad\,\forall p,q$ (8)
 $B_{y}$ $=W_{y}B_{0}W_{y}^{\dagger}\quad\forall y$ (9)
-where $\tau:=\omega^{pq\delta(d,2)/2}$ [39] defined with Kronecker delta $\delta$ and the parameterization $x=dp+q$ is used in Eq.(5).
-The index $x\in(0,d^{2})$ is now expressed with $p,q\in[0,d)$ satisfying $(p,q)\neq(0,0)$ and $y\in[0,d^{2})$.
-Alice’s observables in (5) are given as $d^{2}-1$ WH operators.
-Bob’s observables are generated from $B_{0}$ as Eq.(6).
-An important part of our method is to construct $B_{0}$ such that it has the fiducial state as its eigenstate.
+where $\tau:=\omega^{pq\delta(d,2)/2}$ [47] is a constant defined with Kronecker delta, $\delta$.
+The parameterization $x(p,q):=dp+q$ is used in Eq.(8).
+It allows one to express Alice’s measurements with a single measurement index, $x$.
+Measurements are specified with $x\in(0,d^{2})$ or $p,q\in[0,d)$ satisfying $(p,q)\neq(0,0)$ and $y\in[0,d^{2})$.
+Alice’s measurements in (8) are given as $d^{2}-1$ WH operators when $d\geq 3$.
+Bob’s measurements are generated from $B_{0}$ as Eq.(9).
+An important part of our method is to construct $B_{0}$ such that it has at least one fiducial state as its eigenstate.
 This choice of $B_{0}$ makes it possible to realize WH group covariant SICs on Bob’s eigenspace.
-Consider $B_{0}=\sum_{\beta}\omega^{\beta}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|$ where the eigenspace is defined as $\Phi(\varphi):=\{\left|\varphi_{\beta}\right>|\beta\in[0,d)\}$, a set of $d$ orthonormal states including the fiducial state $\left|\varphi\right>$.
-Without the loss of generality, one can let $\left|\varphi\right>:=\left|\varphi_{d-1}\right>$.
+Consider $B_{0}=\sum_{\beta}\omega^{\beta}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|$ whose eigenspace is defined as $\Phi_{d}:=\{\left|\varphi_{\beta}\right>|\beta\in[0,d)\}$, a set of $d$ orthonormal states including the fiducial state $\left|\varphi\right>$.
+Without loss of generality, one can let $\left|\varphi\right>:=\left|\varphi_{d-1}\right>$.
 Then we have,
 $B_{y}$ $=\sum_{\beta}\omega^{\beta}W_{y}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|W_{y}^{\dagger}$ (10)
-where the eigenbases of $B_{y}$ are now $\{W_{y}\left|\varphi_{\beta}\right>|\beta\in[0,d)\}$.
-Consequently, the set of $(d-1)$-th eigenstates of all the Bob’s observables comprises the $d^{2}$ WH group covariant SICs, $\{W_{y}\left|\varphi\right>|y\in[0,d^{2})\}$.
+where $\left|\varphi\right>:=\left|\varphi_{d-1}\right>$.
+The above spectral decomposition shows that the eigenbases of $B_{y}$ are now $\{W_{y}\left|\varphi_{\beta}\right>|\beta\in[0,d)\}$.
+Consequently, as $W_{y}\left|\varphi_{d-1}\right>=W_{y}\left|\varphi\right>$, collecting the set of $(d-1)$-th eigenstates of all $d^{2}$ Bob’s measurements, WH group covariant SICs $\{W_{y}\left|\varphi\right>|y\in[0,d^{2})\}$ is obtained.
+Note that conditions (ii) and (iii) are systematically satisfied with the choice (8) and (10).
 […]
-$\tilde{Q}:=\frac{1}{2}\left[\sum_{n}\lVert F_{n}\rVert^{2}+d^{2}(d-1)\right]$ (8)
-when $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ for all cases of $a\neq x$.
-Second, we solve the saturation condition with optimal realization, $\left|\phi_{d}^{+}\right>$ and observables (5) and (7).
-Saturation condition is given as $(B_{y}^{n})^{*}=D_{y}^{n}=\sum_{x}f_{x,y}^{n}W_{x}^{n}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ where $\mathbf{A}_{n}:=\tau^{n}(W_{1}^{n},W_{2}^{n},\ldots,W_{d^{2}-1}^{n})^{\intercal}$.
-The solution $\mathbf{c}_{y}^{n}$ completely derives the coefficient matrices we after.
-In sum, we have to solve $(B_{y}^{n})^{*}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ about $\{f_{x,y}^{n}\}$ under constraint $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ $\forall a\neq x$.
-The solution is derived in the below theorem.
-Theorem. Consider prime $d$, for which WH group covariant SICs is defined. Then, there always exists the expression $S$ defined with,
+$\tilde{Q}(d):=\frac{1}{2}\left[\sum_{n}\lVert F_{n}\rVert^{2}+d^{2}(d-1)\right]$ (11)
+when $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ for $a\neq x$.
+Second, we solve the saturation condition with generalized EBI correlation, specifically correlation from $\left|\phi_{d}^{+}\right>$ and measurements (8) and (10).
+As previously mentioned, the condition can be satisfied with the maximally entangled state and $(B_{y}^{n})^{*}=D_{y}^{n}=\sum_{x}f_{x,y}^{n}A_{x}^{n}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ where $\mathbf{A}_{n}:=(A_{1}^{n},A_{2}^{n},\ldots,A_{d^{2}-1}^{n})^{\intercal}$.
+From the solution $\mathbf{c}_{y}^{n}$, one obtain a correlation satisfying the saturation condition.
+In sum, we solve $(B_{y}^{n})^{*}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ about $\mathbf{c}_{y}^{n}$, or about $\{f_{x,y}^{n}\}$, under constraint $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ $\forall a\neq x$.
+The derivation is given in the following theorem.
+Theorem. Consider prime dimension $d$, in which WH group covariant SICs is defined. Then, there always exists the expression $S$ defined with
 $f_{dp+q,dr+s}^{n}=\omega^{-n(ps+qr)}f_{dp+q,0}^{n}$ (12)
-for $p,q,r,s\in[0,d)$, $dp+q\neq 0$ that satisfies $Q=\tilde{Q}$ with generalized EBI correlation.
-Proof.—Consider optimal $B_{0}=\sum_{\beta=0}^{d-1}\omega^{\beta}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|$ defined with eigenspace $\Phi(\varphi)$ containing $\left|\varphi\right>_{d-1}:=\left|\varphi\right>$.
-This choice of $B_{0}$ realizes SICs in the eigenspace as explained.
-Saturation condition is re-written with (6) as
-$(B_{y}^{n})^{*}$ $=(W_{y}B_{0}^{n}W_{y}^{\dagger})^{*}$
-$=\mathbf{c}_{0}^{n}\cdot(W_{y}^{*}\mathbf{A}_{n}W_{y}^{\intercal})$
-$=\sum_{p,q}\omega^{-n(ps+qr)}f_{dp+q,0}^{n}(\mathbf{A}_{n})_{dp+q}$
-$=\sum_{p,q}f_{dp+q,dr+s}^{n}(\mathbf{A}_{n})_{dp+q}$ (10)
-where the notation $(\cdot)_{x}$ stands for the $x$-th element.
-The second equation is derived from saturation condition, $B_{0}^{n}=(\mathbf{c}_{0}^{n}\cdot\mathbf{A}_{n})^{*}$.
-Substituting $(\mathbf{A}_{n})_{dp+q}=(\tau X^{p}Z^{q})^{n}$ and $W_{y}=W_{dr+s}=X^{r}Z^{s}$ into RHS of the second equation and suitably applying the canonical commutation relation, $Z^{q}X^{p}=\omega^{pq}X^{p}Z^{q}$, derives the third equation.
-We used the parameterizations $x=dp+q$ and $y=dr+s$ for the substitution.
-The equation (10) is obtained from $(B_{dr+s}^{n})^{*}=\mathbf{c}_{dr+s}^{n}\cdot\mathbf{A}_{n}$ and it results in Eq.(9).
-Coefficients in Eq.(9) also satisfies the decomposition condition.
-Substituting (9) into $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}$, one can find that it reduces to 0 for all $n$ if $a\neq x$.
-Finally, showing the saturation condition is always solvable completes the theorem.
+for $p,q,r,s\in[0,d)$, $dp+q\neq 0$ satisfying $Q=\tilde{Q}$ with generalized EBI correlation.
+Proof.—Consider an optimal measurement $B_{0}=\sum_{\beta=0}^{d-1}\omega^{\beta}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|$ whose eigenspace contains a fiducial state $\left|\varphi\right>_{d-1}:=\left|\varphi\right>$.
+This choice of $B_{0}$ realizes SICs in the eigenspace, as previously explained.
+Applying saturation condition for $y=0$, $(B_{0}^{n})^{*}=\mathbf{c}_{0}^{n}\cdot\mathbf{A}_{n}$, into the RHS of Eq.(9), it is derived that
+$(W_{y}B_{0}^{n}W_{y}^{\dagger})^{*}$ $=\mathbf{c}_{0}^{n}\cdot(W_{y}^{*}\mathbf{A}_{n}W_{y}^{\intercal})$
+$=\sum_{p,q}\omega^{-n(ps+qr)}f_{dp+q,0}^{n}A_{dp+q}^{n}.$ (13)
+Substituting $A_{dp+q}^{n}=(\tau X^{p}Z^{q})^{n}$ and $W_{y}=W_{dr+s}=X^{r}Z^{s}$ into the RHS of the first equation and suitably applying the canonical commutation relation, $Z^{q}X^{p}=\omega^{pq}X^{p}Z^{q}$, derives Eq.(13).
+The parameterization $x=dp+q$ and $y=dr+s$ are used in the derivation.
+While, the saturation condition for arbitrary $y$, $(B_{y}^{n})^{*}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$, reads
+$(B_{y}^{n})^{*}$ $=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$
+$=\sum_{x}f_{x,y}^{n}A_{x}^{n}$
+$=\sum_{p,q}f_{dp+q,dr+s}^{n}A_{dp+q}^{n}$ (14)
+where the same parameterization $x=dp+q$ and $y=dr+s$ are used for comparison with (13).
+Consequently, one obtain the claimed Eq.(12) from equating (13) and (14).
+The expression given as (12) also satisfies the decomposition condition.
+Substituting (12) into $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}$, one finds that it reduces to 0 for all $n$ if $a\neq x$.
+Showing saturation condition is always solvable completes the proof.
 The elements of $\mathbf{A}_{n}$ is the unitary operator bases for traceless operator [31].
-It means that the traceless operator $(B_{0}^{n})^{*}$ can always be expressed as $\mathbf{c}_{0}^{n}\cdot\mathbf{A}_{n}$ with suitable $\mathbf{c}_{0}^{n}$.
+It means that the traceless operator $(B_{0}^{n})^{*}$ can always be expressed as a linear combination form $\mathbf{c}_{0}^{n}\cdot\mathbf{A}_{n}$ with suitable $\mathbf{c}_{0}^{n}$.
 $\blacksquare$
-The above theorem states that the correlation function maximized by generalized EBI correlation always can be found from Eq.(9) with WH group covariant SICs.
-One has freedom in the choice of SICs, thus $B_{0}$, which determines the coefficients $\{f_{x,0}^{n}\}$ with $(B_{0}^{n})^{*}=\sum_{x}f_{x,0}^{n}(\mathbf{A}_{n})_{x}$.
-One can construct different correlation functions by varying $\{f_{x,0}^{n}\}$ of Eq.(9) with different choices of SICs.
-We also note that establishing $\{A_{x}\}$ as unitary operator bases [34] for traceless operator is advantageous for restricting the number of measurements, as the bases are minimum number of unitary operators that express, in linear combination form, an arbitrary traceless unitary operator, in our case $B_{0}^{n}$.
-IV Derivation of Bell inequality
-IV.1 Gisin’s elegant Bell inequality
+The above theorem shows that the Bell expression maximized by generalized EBI correlation can always be defined with (12).
+It is noteworthy that one has freedom in choice of SICs which determine $f_{x,0}^{n}$, the variable parameters in the RHS of (12).
+With a chosen WH covariant SICs and the corresponding fiducial state, one can construct the optimal $B_{0}$ as previously explained.
+Then the coefficients for $y=0$, $f_{x,0}^{n}$, are derived from the saturation condition, $(B_{0}^{n})^{*}=\sum_{x}f_{x,0}^{n}(\mathbf{A}_{n})_{x}$.
+Different Bell expressions can be generated by varying $f_{x,0}^{n}$ in the RHS of Eq.(12) with different choices of SICs.
+We also note that assuming Alice’s measurements as unitary operator bases [31] for traceless operator is advantageous for restricting the number of measurements in our method.
+The traceless unitary operator bases $\{U_{x}\}$ are the minimum number of unitary operators that satisfy $V=\sum_{x}v_{x}U_{x}$ for an arbitrary traceless unitary operator $V$.
+Therefore, with our choice of Alice’s measurements, $B_{0}$ constructed from any fiducial vector can be considered to satisfy the saturation condition, $(B_{0}^{n})^{*}=\sum_{x}f_{x,0}^{n}(\mathbf{A}_{n})_{x}$.
+In other words, any WH covariant SICs can be considered.
+As a Bell test requires random inputs to close freedom-of-choice loophole [50], reducing the number of settings can be advantageous for lowering input randomness.
+Also, fewer measurements might beneficial to restrict the number of experimental parameters to be controlled.
+5 Derivation of Bell inequality
+5.1 Gisin’s elegant Bell inequality
 […]
-IV.2 Bell inequality for two-qutrit system
+5.2 Bell inequality for two-qutrit system
 We derive a Bell inequality maximally violated by generalized EBI correlation in $d=3$.
 For $d=3$, there are infinitely many fiducial states generating WH covariant SICs, for example $\left|\psi\right>=1/\sqrt{2}(0,1,-\omega^{\theta})$ for $20\leq\theta\leq 1/$ 2 [49].
-We choose the one with $\theta=0$, $\left|\varphi\right>:=1/\sqrt{2}(0,1,-1)=1/\sqrt{2}\left(\left|1\right>-\left|2\right>\right)$ which is known to generate the Hesse SICs [42].
-With the set of orthonormal bases $\left|\varphi_{0}\right>:=\left|0\right>$, $\left|\varphi_{1}\right>:=1/\sqrt{2}(\left|1\right>+\left|2\right>)$, and $\left|\varphi_{2}\right>:=\left|\varphi\right>$, optimal $B_{0}$ is
+We choose one with $\theta=0$, $\left|\varphi\right>=1/\sqrt{2}(0,1,-1)=1/\sqrt{2}\left(\left|1\right>-\left|2\right>\right)$ which is known to generate the Hesse SICs [42].
+With the set of orthonormal bases $\left|\varphi_{0}\right>:=\left|0\right>$, $\left|\varphi_{1}\right>:=1/\sqrt{2}(\left|1\right>+\left|2\right>)$, and $\left|\varphi_{2}\right>:=\left|\varphi\right>$, optimal $B_{0}$ is constructed as
 $B_{0}=\sum_{\beta}\omega^{\beta}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|=\begin{pmatrix}1&0&0\\ 0&-1/2&i\sqrt{3}/2\\ 0&i\sqrt{3}/2&-1/2\end{pmatrix}.$ (15)
-It is worth mentioning here that our choice of $\left|\varphi_{1}\right>$ is also a fiducial state generating WH covariant SICs.
-By solving the saturation condition of SOS decomposition, $(B_{0})^{*}=\mathbf{c}_{0}^{1}\cdot\mathbf{A}_{1}$, the first column of $F_{1}$ is easily obtained as $\mathbf{c}_{0}^{1}=(f_{1,0}^{1},f_{2,0}^{1},\ldots,f_{8,0}^{1})^{\intercal}=(1/2)(1,1,\lambda,\nu,\mu,\lambda,\mu,\nu)^{\intercal}$ with $\lambda:=-i/\sqrt{3}$, $\mu:=\omega\lambda$ and $\nu:=\omega^{2}\lambda$.
-Substituting $\mathbf{c}_{0}$ into Eq.(9) and using the real-valuedness condition $F_{2}=F_{1}^{*}$, the coefficients are fully generated as in [40].
-The derived coefficients can be further simplified in a functional form, $f_{dp+q,0}^{n}=1/2$ for $p=0$ and $f_{dp+q,0}^{n}=\omega^{-n(pq+3/2)+3/4}/(2\sqrt{3})$ for $p\neq 0$.
-Substituting the derived coefficients into (2), a Bell inequality maximally violated by generalized EBI correlation in two-qurtrit system is derived as,
-$S_{3}:=\sum_{n=1}^{2}\sum_{\begin{subarray}{c}p,q,r,s=0\\ 3p+q\neq 0\end{subarray}}^{2}\omega^{-n(ps+qr)}f_{3p+q,0}^{n}\left\langle A_{3p+q}^{n}B_{3r+s}^{n}\right\rangle\leq 15.$ (12)
-The upper bound $L=15$ can be obtained in the set of local deterministic correlations satisfying $P(\alpha\beta|xy)=1$ for certain combination of $\alpha,\beta$ and otherwise the probability is 0.
-The optimal strategy is found as, for all $x$, assigning $P(00|xy)$, $P(01|xy)$, $P(02|xy)$ as 1 respectively for $y=0,1,2$, $y=3,6$, and $y=4,5,7,8$.
-Violation is obtained from the maximal quantum value, $\tilde{Q}=18$, evaluated from (8) with $\lVert F_{n}\rVert=9$ for $n=1,2$.
-The upper bound is attained by two parties performing measurements, $A_{x}=W_{x}$ and $B_{y}=W_{y}B_{0}W_{y}^{\dagger}$, on the two-qutrit maximally entangled state when $x\in(0,9)$ and $y\in[0,9)$.
-As expected, $2d^{2}$ eigenstates, $W_{y}\left|\varphi_{1}\right>$ and $W_{y}\left|\varphi_{2}\right>$ for all $y$, comprise dual sets of $d^{2}$ SICs and eigenstates of four optimal observables for $x=1,3,4,5$, which are given as $Z,X,XZ,XZ^{2}$, defines the complete sets of MUBs [40].
-We remark that the Bell inequality defined in (12) shows stronger violation than known Bell inequalities of similar classes.
-The critical visibility $\nu^{c}$, the smallest visibility $\nu$ of the isotropic state $\rho=\nu\left|\phi_{d}^{+}\right>\left<\phi_{d}^{+}\right|+(1-\nu)\mathds{I}/d^{2}$ to show the violation, is 83.33% for $S_{3}$.
-It is smaller than that of Bell inequalities including ones defined in $d=3$ for MUBs, 96.77% [19], 96.63% [20], for SICs, 96.41% [20], and original EBI, 86.6%.
+It is worth mentioning here that our choice of $\left|\varphi_{1}\right>$ is also a fiducial state for another WH covariant SICs.
+By solving the saturation condition, $(B_{0})^{*}=\mathbf{c}_{0}^{1}\cdot\mathbf{A}_{1}$, the first column of $F_{1}$ is straightforwardly obtained as $\mathbf{c}_{0}^{1}=(f_{1,0}^{1},f_{2,0}^{1},\ldots,f_{8,0}^{1})^{\intercal}=(1/2)(1,1,\lambda,\nu,\mu,\lambda,\mu,\nu)^{\intercal}$ with $\lambda:=-i/\sqrt{3}$, $\mu:=\omega\lambda$ and $\nu:=\omega^{2}\lambda$.
+Substituting $\mathbf{c}_{0}$ into Eq.(12) and using the real-valuedness condition $F_{2}=F_{1}^{*}$, the coefficients are fully generated as presented in [48].
+The derived coefficients for the first column can be further simplified in a functional form,
+$f_{dp+q,0}^{n}=\begin{cases}\frac{1}{2},\quad(p=0)\\ \frac{i}{2\sqrt{3}}(-1)^{n}\omega^{-npq},\quad(p\neq 0).\end{cases}$ (16)
+Substituting the derived coefficients (16) into (12), a Bell expression is defined as:
+$S_{3}:=\sum_{n=1}^{2}\sum_{\begin{subarray}{c}p,q,r,s=0\\ 3p+q\neq 0\end{subarray}}^{2}\omega^{-n(ps+qr)}f_{3p+q,0}^{n}\left\langle A_{3p+q}^{n}B_{3r+s}^{n}\right\rangle.$ (17)
+A Bell inequality for generalized EBI correlation in two-qutrit system is derived as $S_{3}\leq 15$.
+The upper bound $L=15$ can be obtained in the set of local deterministic correlations obeying $P(\alpha\beta|xy)=1$ for a certain combination of $\alpha,\beta$ and otherwise the probability is 0.
+The optimal strategy is found as, for any $x$ and respectively for $y=0,1,2$, $y=3,6$, and $y=4,5,7,8$, assigning $P(00|xy)$, $P(01|xy)$, $P(02|xy)$ as 1.
+Violation is obtained from the maximal quantum value, $\tilde{Q}=18$, evaluated from (11) with $\lVert F_{1}\rVert=\lVert F_{2}\rVert=9$.
+It is obtained by two parties performing measurements, $A_{x}=W_{x}$ and $B_{y}=W_{y}B_{0}W_{y}^{\dagger}$, on a two-qutrit maximally entangled state $|\phi^{+}_{3}\rangle$ when $x\in(0,9)$ and $y\in[0,9)$.
+As we considered two fiducial states, $\left|\varphi_{1}\right>$ and $\left|\varphi_{2}\right>$, $2d^{2}$ eigenstates on Bob’s side $\{W_{y}\left|\varphi_{1}\right>|y\in[0,9)\}$ and $\{W_{y}\left|\varphi_{2}\right>|y\in[0,9)\}$ make dual SICs.
+The eigenstates of Alice’s four optimal observables for $x=1,3,4,5$, namely $Z,X,XZ,XZ^{2}$, define the complete MUBs [48].
+We remark that the Bell inequality $S_{3}$ shows stronger violation than known Bell inequalities of similar classes.
+The critical visibility $\nu^{c}$, the smallest visibility $\nu$ of the isotropic state $\rho=\nu|\phi_{d}^{+}\rangle\langle\phi_{d}^{+}|+(1-\nu)\mathds{I}/d^{2}$ to show the violation, is 83.33% for $S_{3}$.
+It is smaller than that of Bell inequalities defined for $d=3$ MUBs, 96.77% [20], 96.63% [21], SICs, 96.41% [21], and original EBI, 86.6%.
 In addition, the number of settings required for Alice is 8 for $S_{3}$ and it is smaller than the case of known Bell inequality for $d=3$ SICs which requires $\binom{d^{2}}{2}=36$ settings for Alice when the number of Bob’s setting is same as 9 [21].
-V Conclusion
-We presented a method to construct correlation function for which the quantum maximum is always obtained in a simple analytic form from the generalized EBI correlation.
-Bell inequality for two-qutrit system derived in our method shows larger violation compared to known Bell inequalities of similar classes, including original EBI.
-Our work lays the foundation for the future studies for self-testing MUBs and SICs in high dimensional system and DI quantum cryptographic protocols based on it.
-Acknowledgements.
+6 Conclusion
+We present a method to construct Bell expressions for which the quantum maximum is always obtained in a simple analytic form from the generalized elegant Bell inequality (EBI) correlation.
+That any WH SICs can be considered for the construction provides versatile nature for our scheme.
+It is also discussed that the number of measurement settings are restricted while maintaining the versatility.
+For example, we derived the original EBI and a novel Bell inequality for generalized EBI correlation by applying our method respectively for two and three dimension.
+It is notable that the derived Bell inequality shows larger violation compared to known Bell inequalities, either for three dimensional MUBs or SICs.
+It improves the visibility of isotropic state more than 13%.
+Our primary contribution lies in suggesting a method to derive Bell expressions maximized by generalized EBI correlation.
+The exploration of Bell inequalities in arbitrary higher dimensions represents an interesting direction for future research, requiring broader analysis such as finding suitable choice of SICs in high dimension.
+We believe our work provides a key framework for such exploration.
+This work also naturally establishes a foundation for future research on self-testing MUBs and SICs in high-dimensional systems, as well as DI quantum cryptographic protocols which would be built upon them.
 This research was supported by Korea Institute of Science and Technology Information(KISTI).(No.
-(KISTI)K24L4M1C2) and the National Research Council of Science & Technology(NST) grant by the Korea government (MSIT) (No.
+(KISTI)K25L5M2C2) and the National Research Council of Science & Technology(NST) grant by the Korea government (MSIT) (No.
 CAP22053-000).
@@
 NRF-2022M3K2A1083890).
-Appendix A Symmetries on the coefficient matrices and optimal measurements
-We present the coefficient matrix for $S_{3}$ and MUBs and SICs that gives the optimal violation.
-Although the symmetries in $S_{3}$ can be expressed with an analytic equation in the main text, the explicit form of the coefficient matrix, MUBs and SICs would show the symmetries in the forms that can be seen at a glance.
-The coefficient matrices are given as $8\times 9$ matrices satisfying real-valuedness condition, $F_{2}=F_{1}^{*}$ as follows
+Appendix A MUBs in Alice’s eigenspace
+Complete MUBs are realized in the eigenspace of Alice’s $d^{2}-1$ measurements under condition (ii) of generalized EBI correlation.
+As previously noted in section 3, it is straightforwardly shown from $\mathcal{C}:=\{Z,X,XZ,\ldots,XZ^{d-1}\}\subset\mathcal{W}_{d}$ and the fact that the eigenstates of the $d+1$ unitary operators in $\mathcal{C}$ are complete MUBs for prime dimension [43].
+We prove another notable property regarding MUBs on Alice’s eigenspace in the following proposition.
+Proposition. The set of all eigenstates of Alice’s measurements defined in generalized EBI correlation are $(d-1)$-fold complete MUBs.
+Proof.
+The measurements generating complete MUBs, $\mathcal{C}$, are divided into two subsets $\mathcal{C}_{1}:=\{Z\}$ and $\mathcal{C}_{2}:=\{XZ^{q}|q\in[0,d)\}$ such that $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$.
+The set of Alice’s measurements $\mathcal{A}:=\{X^{p}Z^{q}|(p,q)\neq(0,0)\}$ can be similarly divided with $\mathcal{A}_{1}:=\{Z^{q}|q\neq 0\}$ and $\mathcal{A}_{2}:=\{X^{p}Z^{q}|p\neq 0\}$, so that $\mathcal{A}=\mathcal{A}_{1}\cup\mathcal{A}_{2}$.
+Then the above proposition is proved by
+$V(\mathcal{C}_{1})$ $=V(\mathcal{A}_{1}),\,\mbox{and}$ (18)
+$V(\mathcal{C}_{2})$ $=V(\mathcal{A}_{2})$ (19)
+where $V(O)$ denotes a set of unit eigenvectors of all elements in a given set, $O$. The Eq.(18) are easily proved, as taking the power of a unitary matrix preserves the eigenspace.
+Applying the same property of unitary matrices to the left-hand side (LHS) of the Eq.(19) derives $V(\mathcal{C}_{2})=V(\{(XZ^{q})^{p}|p\neq 0\})$.
+It also holds that $(XZ^{q})^{p}=\omega^{n}X^{p}Z^{\dot{(pq)}}$ from the canonical commutation relation, $Z^{q}X^{p}=\omega^{pq}X^{p}Z^{q}$, where $n(p,q):=qp(p-1)/2$ is an integer and the overdot denotes the smallest non-negative $d$ modulo value.
+Then, one obtains $V(\mathcal{C}_{2})=V(\{X^{p}Z^{\dot{(pq)}}|p\neq 0\})$.
+Consequently, the Eq.(19) can be converted into
+$V(\{X^{p}Z^{\dot{(pq)}}|p\neq 0\})=V(\mathcal{A}_{2})=V(\{X^{p}Z^{q}|p\neq 0\}).$ (20)
+This equation can be proved with the below equivalence
+$q_{1}\neq q_{2}\Leftrightarrow\dot{(pq_{1})}\neq\dot{(pq_{2})}$ (21)
+for $p\in(0,d)$ and $q_{1},q_{2}\in[0,d)$.
+It means that all different values of $q$ ($q_{1}$ and $q_{2}$) implies different values of $\dot{(pq_{1})}$ and $\dot{(pq_{2})}$.
+The situation results in the one-to-one correspondence between set of $q$ and set of $\dot{(pq)}$ for any given non-zero $p$.
+This is sufficient to proves Eq.(20).
+We prove the rightward relation ($\Rightarrow$) first.
+The least non-negative $d$-modulo value of any non-negative integer $x$ can be expressed as $\dot{x}=x-d\lfloor x\rfloor$ with the quotient denoted by $\lfloor x\rfloor$.
+Using this expression, the RHS of (21) can be modified as
+$\frac{p(q_{1}-q_{2})}{d}\neq\left\lfloor\frac{pq_{2}}{d}\right\rfloor-\left\lfloor\frac{pq_{1}}{d}\right\rfloor.$ (22)
+Now suppose that the LHS of (21), $q_{1}\neq q_{2}$ Without loss of generality one can consider the case $q_{1}>q_{2}$.
+Then, as $d$ is prime and $p$ and $q_{1}-q_{2}$ are all smaller than $d$, the LHS of (22) is always a fractional number.
+While the RHS clearly is an integer.
+It proves (22), thus rightward relation.
+The leftward relation ($\Leftarrow$) is easily proved from its contraposition, $q_{1}=q_{2}\Rightarrow\dot{(pq_{1})}=\dot{(pq_{2})}$.
+$\blacksquare$
+We note that the above proposition implies that any one of the eigenstates of Alice’s measurements fall into one of the eigenstates of $Z,X,XZ,\ldots,XZ^{d-1}$ that comprise the complete MUBs.
+Appendix B Quantum bound
+Detailed version of the derivation of quantum upper bound $\tilde{Q}$ is given in this section.
+As shown in the main text, the SOS decomposition ¯$\bar{\mathcal{B}}=\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}\geq 0$ can be modified to
+$\mathcal{B}\leq\frac{1}{2}\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}\otimes\mathds{1}+\frac{1}{2}\sum_{n,y}\mathds{1}\otimes B_{y}^{n}(B_{y}^{n})^{\dagger}$ (23)
+by substituting $P_{y}^{n}:=D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}$.
+The second term of the RHS of (23) is reduced to $d^{2}(d-1)/2\mathds{I}$ from the unitarity of $B_{y}^{n}$.
+The first term of the RHS is converted into an identity matrix multiplied by some real constant by substituting $D_{y}^{n}:=\sum_{x}f_{x,y}^{n}A_{x}^{n}$ into $\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}$, as follows,
+$\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}$ $=\sum_{n,y}\left(\sum_{a}f_{a,y}^{n}A_{a}^{n}\right)^{\dagger}\sum_{x}f_{x,y}^{n}A_{x}^{n}$
+$=\sum_{n,a,x}\left[\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}\right](A_{a}^{n})^{\dagger}A_{x}^{n}$
+$=\sum_{n,x,y}|f_{x,y}^{n}|^{2}\mathds{1}+\sum_{n}\sum_{a\neq x}\left[\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}\right](A_{a}^{n})^{\dagger}A_{x}^{n}$
+$=\sum_{n}\lVert F_{n}\rVert^{2}\mathds{1}+\sum_{n}\sum_{a\neq x}(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}(A_{a}^{n})^{\dagger}A_{x}^{n}.$ (24)
+[…]
+$\tilde{Q}:=\frac{1}{2}\left[\sum_{n}\lVert F_{n}\rVert^{2}+d^{2}(d-1)\right].$ (25)
+Appendix C Coefficient matrix, MUBs and SICs in three dimension
+We present the coefficient matrix for $S_{3}$, and MUBs and SICs that gives the optimal violation.
+Although the symmetries in $S_{3}$ can be expressed with the analytic equation in the main text, the explicit forms of the coefficient matrix, MUBs and SICs would reveal the symmetries such that it can be seen at a glance.
+The coefficient matrices are given as $8\times 9$ matrices as follows:
 $F_{1}:=\frac{1}{2}$ $\begin{bmatrix}1&1&1&\omega^{2}&\omega^{2}&\omega^{2}&\omega&\omega&\omega\\ 1&1&1&\omega&\omega&\omega&\omega^{2}&\omega^{2}&\omega^{2}\\ \lambda&\nu&\mu&\lambda&\nu&\mu&\lambda&\nu&\mu\\ \nu&\mu&\lambda&\mu&\lambda&\nu&\lambda&\nu&\mu\\ \mu&\lambda&\nu&\nu&\mu&\lambda&\lambda&\nu&\mu\\ \lambda&\mu&\nu&\lambda&\mu&\nu&\lambda&\mu&\nu\\ \mu&\nu&\lambda&\lambda&\mu&\nu&\nu&\lambda&\mu\\ \nu&\lambda&\mu&\lambda&\mu&\nu&\mu&\nu&\lambda\end{bmatrix}$
-when $\lambda:=-i/\sqrt{3}$, $\mu:=\omega\lambda$ and $\nu:=\omega^{2}\lambda$. One can find the first column generates the others following, $f_{dp+q,dr+s}^{1}=\omega^{-(ps+qr)}f_{3p+q,0}^{1}$.
+when $\lambda:=-i/\sqrt{3}$, $\mu:=\omega\lambda$ and $\nu:=\omega^{2}\lambda$.
+As expected, one can find that Eq.(12) holds.
+All the other coefficients is obtained from the realvaluedness condition, $F_{2}=F_{1}^{*}$.
 […]
@@
 $\mbox{SIC}_{2}$ $=\frac{1}{\sqrt{2}}\begin{bmatrix}0&-1&1&0&-\omega^{2}&\omega&0&-\omega&\omega^{2}\\ 1&0&-1&\omega&0&-\omega^{2}&\omega^{2}&0&-\omega\\ -1&1&0&-\omega^{2}&\omega&0&-\omega&\omega^{2}&0\end{bmatrix}$
-where each column of $\mbox{SIC}_{1}$ and $\mbox{SIC}_{2}$ defines SICs generated from $\left|\varphi_{1}\right>$ and $\left|\varphi_{2}\right>$ respectively.
+where the columns of $\mbox{SIC}_{\beta}$ correspond to the normalized vectors defining 9 SIC-POVM elements generated from $\left|\varphi_{\beta}\right>$.
 On the other side, the eigenstates of Alice’s optimal observables $A_{1}=Z$, $A_{3}=X$, $A_{4}=XZ$, $A_{5}=XZ^{2}$ respectively define the complete mutually unbiased bases,
@@
 $\mbox{MUB}_{3}=\frac{1}{\sqrt{3}}\begin{bmatrix}\omega^{2}&\omega&1\\ \omega^{2}&1&\omega\\ 1&1&1\end{bmatrix},\,\mbox{MUB}_{4}=\frac{1}{\sqrt{3}}\begin{bmatrix}\omega&1&\omega^{2}\\ \omega&\omega^{2}&1\\ 1&1&1\end{bmatrix}.$
-We note that the eigenvectors of other settings $A_{2},A_{6},A_{7},A_{8}$ also define the above four MUBs.
-It can be easily shown from the fact that they are unitary and up to phase equivalent to $A_{1}^{2},A_{3}^{2},A_{5}^{2},A_{4}^{2}$ respectively.
-Appendix B Derivation of the quantum bound
-Detailed version of the derivation of quantum upper bound, $\tilde{Q}$, is presented in this section.
-As shown in the main text, the SOS decomposition $\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}\geq 0$ can be modified as
-$\mathcal{B}\leq\frac{1}{2}\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}\otimes\mathds{1}+\frac{1}{2}\sum_{n,y}\mathds{1}\otimes B_{y}^{n}(B_{y}^{n})^{\dagger}$ (13)
-by substituting $P_{y}^{n}:=D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}$.
-The second term of the RHS of (13) is reduced to $d^{2}(d-1)/2\mathds{1}$ from the unitarity of $B_{y}^{n}$.
-The first term of the RHS is derived as the desirable form by substituting $D_{y}^{n}:=\sum_{x}f_{x,y}^{n}A_{x}^{n}$ into $\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}$, as follows,
-$\sum_{n,y}\left(\sum_{a}f_{a,y}^{n}A_{a}^{n}\right)^{\dagger}\sum_{x}f_{x,y}^{n}A_{x}^{n}$
-$=\sum_{n,a,x}\left[\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}\right](A_{a}^{n})^{\dagger}A_{x}^{n}$
-$=\sum_{n,x,y}|f_{x,y}^{n}|^{2}\mathds{1}+\sum_{n}\sum_{a\neq x}\left[\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}\right](A_{a}^{n})^{\dagger}A_{x}^{n}$
-$=\sum_{n}\lVert F_{n}\rVert^{2}\mathds{1}+\sum_{a\neq x}(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}(A_{a}^{n})^{\dagger}A_{x}^{n}$ (14)
-where the Frobeniuns norm $\lVert F_{n}\rVert:=(\sum_{x,y}|f_{x,y}^{n}|^{2})^{1/2}$. Consequently, the upper bound in (13) is reduced to
-$\tilde{Q}:=\frac{1}{2}\left[\sum_{n}\lVert F_{n}\rVert^{2}+d^{2}(d-1)\right]$ (15)
-under conditions $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ for all $n,a\neq x$.
+As previously discussed in Appendix A, the eigenvectors of other settings $A_{2},A_{6},A_{7},A_{8}$ also define the above four MUBs.